### v2

# AffectGPT-R1: Leveraging Reinforcement Learning for Open-Vocabulary Multimodal Emotion Recognition

Abstract
Open-Vocabulary Multimodal Emotion Recognition (OV-MER) aims to predict emotions without being constrained by predefined label spaces, enabling fine-grained emotion understanding. Unlike traditional discriminative methods, OV-MER leverages generative models, such as large language models, to capture the full spectrum of emotions and employs emotion wheels (EWs) for metric calculation. Previous approaches (e.g., AffectGPT) primarily rely on token-level loss during training. However, this objective is misaligned with the metrics used in OV-MER, while these metrics cannot be optimized via gradient backpropagation. In this paper, we propose AffectGPT-R1, a reinforcement learning framework that formulates EW-based metrics as a reward function and employs a policy-based optimization strategy to maximize this reward. Additionally, we introduce an extra reasoning process and investigate its necessity in OV-MER. To further refine model behavior, we incorporate auxiliary rewards that constrain both reasoning and emotion prediction. To prevent reward hacking, we propose to incorporate length penalties during training. Experimental results show that AffectGPT-R1 achieves substantial improvements on OV-MER. Beyond this task, our approach also enhances generalized emotion understanding, attaining state-of-the-art performance on MER-UniBench. To the best of our knowledge, this is the first work to adapt the R1-style methodology for emotion understanding, revealing the impact of reasoning processes and reinforcement learning in this domain. Our code is provided in the supplementary material and will be released to facilitate future research.
1 Introduction
Emotion is highly related to human cognition, decision-making, and behavior, playing a vital role in our daily life [picard2000affective, xie2024emovit]. Naturally, humans express emotions through multiple modalities, including (micro)-gestures, (micro)-expressions, audio tones, linguistic content, and other cues [ben2021video, chen2023smg, el2011survey, lian2024merbench]. To better understand human emotions, it is essential to integrate these clues, enabling models to capture emotional nuances. This has led to the development of Multimodal Emotion Recognition (MER) [lian2023mer, zhang2025moda]. Traditional MER approaches primarily rely on Ekman’s theory [ekman1992argument, li2023decoupled], which categorizes human emotions into six basic labels: anger, disgust, happiness, sadness, fear, and surprise. However, human emotions extend far beyond these basic categories, and such a limited label space inevitably results in inaccurate emotion descriptions [plutchik1980general, demszky2020goemotions].
Recently, open-vocabulary MER (OV-MER) has emerged as a promising research direction, aiming to shift emotion recognition from constrained categories to the full spectrum of emotion, thereby enabling more nuanced emotion representations [lian2025ov, lian2025affectgpt]. To support this paradigm shift, OV-MER introduces new solutions and metrics. For solutions, it transitions from discriminative to generative approaches, leveraging the extensive vocabulary of large language models (LLMs) to expand the recognition scope of emotion labels. For metrics, it employs emotion wheels (EWs) to capture semantic relationships between distinct emotion words. Appendix A provides a detailed illustration of these metrics. To better address this task, current approaches primarily use token-level loss to align predicted and ground-truth labels [lian2025affectgpt, cheng2024emotion]. However, these methods suffer from a critical misalignment: token-level loss shows limited correlation with EW-based metrics. For instance, sadness and madness exhibit high token-level similarity but low semantic similarity in their underlying emotions. Furthermore, EW-based metrics cannot be directly optimized via gradient backpropagation, introducing significant challenges and limitations to existing solutions.
In this paper, we propose AffectGPT-R1, which leverages EW-based metrics as the reward function and employs reinforcement learning to maximize this reward, enabling the direct optimization of these metrics. Meanwhile, we introduce an additional reasoning process and investigate its necessity in addressing OV-MER. To refine model behaviors, we incorporate auxiliary reward functions to constrain the model’s reasoning and emotion predictions. Through extensive experiments, we analyze the impact of different reward functions and identify the most effective ones. Additionally, during inference, the reward hacking problem [everitt2017reinforcement] arises, where the model tends to generate lengthy predictions with redundant emotional expressions (e.g., happy and happily). To mitigate this, we incorporate length penalties into the rewards, ensuring reward maximization while promoting more concise outputs. Experimental results demonstrate that AffectGPT-R1 achieves significant performance improvements on OV-MER. Our framework also attains state-of-the-art performance on MER-UniBench [lian2025affectgpt], a unified benchmark for generalized emotion understanding, further validating its effectiveness. To the best of our knowledge, this is the first work to apply R1-style training to OV-MER, demonstrating the effect of reasoning processes and reinforcement learning in this domain. The main contributions of this paper are summarized as follows:
  • (Framework) We introduce AffectGPT-R1, the first work to apply an R1-style training strategy to OV-MER, shedding light on the role of reinforcement learning and reasoning processes in this task.
  • (Reward) We systematically investigate various reward functions and conduct an in-depth analysis of their effectiveness, identifying the optimal rewards for OV-MER. To mitigate the reward hacking problem, we introduce length penalties, ensuring the model maximizes rewards while using the fewest emotional words.
  • (Performance) We achieve state-of-the-art performance on OV-MER and MER-UniBench, proving the potential of reinforcement learning for emotion understanding.
2 Related Works
2.1 Emotion Representation
Emotion representation defines the recognition scope of developed models and serves as the foundation of MER [gunes2011emotion]. There are two primary representation methods: dimensional emotion theory [schlosberg1954three] and basic emotion theory [ekman1992argument].
Dimensional Emotion Theory.
This theory uses continuous dimensions to capture the core attributes underlying emotions. For example, Russell [russell1980circumplex] proposed a two-dimensional model, characterizing emotions based on valence (the pleasantness of an emotion) and arousal (the activation level of an emotion). In addition to this, some researchers introduced an additional dominance dimension to reflect the perceived control over emotions [russell1977evidence]. However, these dimensions are abstract and less intuitive for the general public, posing challenges for applications.
Basic Emotion Theory.
This theory typically uses six or eight basic labels to describe emotional states. For example, Ekman [ekman1992argument] argued that human emotions can be categorized into six basic labels, while Plutchik [plutchik1980general] proposed eight primary emotions that serve as the foundation for all others. Nevertheless, human emotions are rich and diverse, extending far beyond these basic labels. Mapping such nuanced emotional expressions into a restricted set of basic labels inevitably leads to the loss of subtle emotional distinctions.
Open-vocabulary Emotion.
To address these limitations, researchers propose an open-vocabulary representation manner [lian2025ov], which removes restrictions on the label space and enables models to predict any emotion. Unlike dimensional emotion theory, it focuses on more intuitive emotion words; Unlike basic emotion theory, it avoids restrictive label spaces, enabling more nuanced emotion representation. This paper focuses on open-vocabulary emotions and introduces AffectGPT-R1, a pioneering work that leverages the R1-style training strategy to tackle this task.
Figure 1: Overall pipeline of AffectGPT-R1. Our training pipeline consists of two phases: cold start training and reinforcement learning. In the first phase, we adopt the framework of AffectGPT but replace the generated content, including both thinking and answers. In the second phase, we propose five rewards and introduce length penalties to mitigate reward hacking, followed by GRPO for policy updates.
2.2 Reinforcement Learning
Reinforcement learning has advanced the upper-bound performance of LLMs, demonstrating impressive results in tasks such as mathematical verification and code execution [guo2025deepseek, shao2024deepseekmath, cobbe2021training]. This method primarily consists of two key components: policy optimization and reward modeling.
Policy Optimization.
During training, we aim to maximize the reward function through policy updates. Proximal Policy Optimization (PPO) [schulman2017proximal] is a widely used algorithm that leverages a value model for advantage estimation and incorporates token-wise KL penalties to prevent excessive optimization. Direct Preference Optimization (DPO) [rafailov2023direct] focuses on directly optimizing policies to align with human preferences, eliminating the need for an explicit reward model (as required in PPO) and instead relying on an implicit reward model. This approach reduces computational costs and engineering complexity. Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] is also a more efficient alternative to PPO, removing the value model and instead using a group-relative baseline computed from different outputs for the same input to guide optimization. Beyond these algorithms, some efforts have been made to enhance training efficiency and stability [hu2025reinforce++, yu2025dapo]. In this paper, we do not focus on designing novel policy optimization algorithms. Instead, we adopt GRPO for policy updates.
Reward Modeling.
Reward systems can be broadly categorized into rule-based and model-based rewards. The former is suitable for tasks with verifiable outcomes. For example, pass or fail unit tests in code generation [le2022coderl], output language requirements in question-answering tasks [guo2025deepseek], and IoU between predicted bounding boxes and ground-truth annotations in object detection [shen2025vlm]. The latter is mainly used for tasks where rewards are difficult to compute, such as safety, helpfulness, and harmlessness [zhang2024safetybench]. In prior work, how to design reward functions tailored for OV-MER remains unexplored. To this end, we introduce multiple rewards to supervise both thinking and answers, deriving insights into the effective rewards for OV-MER.
3 AffectGPT-R1
Our training pipeline consists of two phases: cold start training and reinforcement learning. In the first phase, we leverage a coarse-grained dataset to establish initial capabilities in emotion understanding and format alignment. In the second phase, we use a fine-grained dataset for reward calculation and policy updates. Figure 1 illustrates the overall training pipeline of AffectGPT-R1.
3.1 Notation Definition
This section introduces some necessary notations. Specifically, we use $v$ to represent the video, which contains multimodal clues (such as visual, acoustic, and lexical information), and $q$ to denote the user message. The output $o$ consists of two components: the thinking $o^{t}$ and the answer $o^{a}$. Specifically, $o^{t}$ and $o^{a}$ correspond to the content enclosed within the <<think>> and <</think>> tags and the <<answer>> and <</answer>> tags, respectively. We denote the ground-truth open-vocabulary labels as $y$. Figure 1 clarifies the meanings of these notations. In this paper, our main goal is to predict $y$ based on the video content $v$.
3.2 Cold Start Training
Human emotions are highly dependent on subtle cues [li2022deep, li2022cas], posing significant challenges for current models to capture all these signals and accurately interpret emotional states [lian2024gpt, lu2024gpt]. To enhance emotion understanding, we employ large-scale coarse-grained descriptive emotion datasets for cold-start training. In this phase, we adopt the same model architecture as AffectGPT [lian2025affectgpt] (see Figure 1). This framework leverages a pre-fusion layer to effectively integrate multimodal cues, emphasizing the inherently multimodal nature of human emotions. Unlike the original AffectGPT, which only outputs emotion words, we generate both thinking (enclosed within <<think>> and <</think>> tags) and answers (enclosed within <<answer>> and <</answer>> tags). Since descriptive emotion datasets already follow a thinking-like format [lian2023explainable], we treat these descriptions as the thinking content. Given our primary focus on OV-MER, we use open-vocabulary emotions as the answer content. Figure 1 illustrates an example of the input and output for cold-start training. This output format unifies descriptive and open-vocabulary emotions into a single task, enabling a single model to leverage their correlation and handle both tasks simultaneously. As demonstrated in Section 5.5, the quantity of cold-start data significantly affects both cold-start training performance and subsequent reinforcement learning. This further emphasizes our rationale for using large-scale datasets in this phase.
3.3 Reinforcement Learning
Reinforcement learning relies on policy optimization and reward design. Rather than focusing on improving optimization strategies, we investigate how to design reward models that can better address OV-MER. In this section, we introduce various rewards to guide thinking and answers.
3.3.1 Policy Optimization
We employ GRPO [shao2024deepseekmath] for policy updates. Unlike PPO [schulman2017proximal], which relies on a critic model for advantage estimation, GRPO eliminates this component and instead uses group scores, significantly reducing memory overhead while improving training efficiency. Specifically, for a given input pair $(v,q)$, the policy model $\pi_{\theta_{\text{old}}}$ samples a group of responses $\mathbf{O}=\{o_{1},\cdots,o_{G}\}$, where $o_{i}$ is the $i$-th response. We then compute the reward for each response as $r_{i}=\mathcal{R}(o_{i}|v,q,y)$, where $\mathcal{R}(\cdot)$ is the reward function defined in Section 3.3.2. To evaluate the relative quality of $\{o_{i}\}_{i=1}^{G}$, we normalize the rewards by computing their mean and standard deviation, deriving the advantage as:
$\hat{A}_{i}=\frac{r_{i}-\text{mean}(\{r_{i}\}_{i=1}^{G})}{\text{std}(\{r_{i}\}_{i=1}^{G})}.$ (1)
During training, GRPO encourages the policy model $\pi_{\theta}$ to generate high-reward responses by minimizing the following objective function:
$\mathcal{J}_{\text{GRPO}}(\theta)=$ $\mathbb{E}_{(v,q)\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|v,q)}$ (2)
$\left[\frac{1}{G}\sum_{i=1}^{G}\left(S_{i}-\beta{D}_{\text{KL}}\left(\pi_{\theta}||\pi_{\text{ref}}\right)\right)\right],$
$S_{i}=\min\left(s_{i}\hat{A}_{i},\text{clip}\left(s_{i},1-\epsilon,1+\epsilon\right)\hat{A}_{i}\right),$ (3)
$s_{i}=\frac{\pi_{\theta}(o_{i}|v,q)}{\pi_{\theta_{\text{old}}}(o_{i}|v,q)},$ (4)
where $\pi_{\theta}$ represents the online policy model and $\pi_{\text{ref}}$ serves as the offline reference model. Similar to PPO, GRPO employs $\epsilon$ as the clipping threshold and incorporates a KL divergence penalty term to regularize the divergence between $\pi_{\theta}$ and $\pi_{\text{ref}}$. The combination of clipping and KL regularization effectively prevents extreme policy updates while maintaining training stability.
During experiments, we observed significant divergence between the learned policy distribution and the initial model distribution, and the restrictive effect of the KL penalty term led to performance degradation. This trade-off between policy exploration and regularization has also been proven in prior work [everitt2017reinforcement]. To address this limitation, we remove the KL regularization term from our algorithm and adopt the following simplified objective function for policy update:
$\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{(v,q)\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|v,q)}\frac{1}{G}\sum_{i=1}^{G}S_{i}.$ (5)
3.3.2 Reward Modeling
Reward functions guide the update direction of policy models. In this work, we propose five reward functions, all designed to evaluate thinking and answers. In Section 5.1, we examine the performance of the single reward and combinations of multiple rewards. Our goal is to identify the most effective reward strategy for OV-MER.
Format Reward.
To ensure structured outputs, we define a binary reward function that evaluates whether the model’s response contains both thinking and answer components, returning 1 if compliant and 0 otherwise.
[…]
Accuracy Reward.
We employ the official metrics of OV-MER to calculate accuracy rewards [lian2025ov]. Unlike discriminative MER, which operates within a fixed label space, OV-MER does not constrain the label space. Therefore, models are permitted to generate synonyms, i.e., emotions with the same meaning but expressed differently, which introduces challenges in evaluation. To mitigate the impact of synonyms, OV-MER introduces EW-based metrics, leveraging the structural information of emotion wheels for clustering. In these wheels, the inner section represents core emotions, while the outer section captures more nuanced emotions. Calculation details can be found in Appendix A:
$\mathcal{R}_{\text{accuracy}}(o|v,q,y)=\text{EW}\left(o^{a},y\right).$ (7)
Alignment Reward.
The accuracy reward ensures the correctness of $o^{a}$. However, without constraints on the thinking process, the generated output may contain unrelated $o^{a}$ and $o^{t}$. To this end, we introduce the alignment reward, ensuring that the reasoning process aligns with the final answer. Specifically, for each output, we extract emotion words $e^{t}$ from $o^{t}$ and compute their similarity with $o^{a}$:
$\mathcal{R}_{\text{alignment}}(o|v,q,y)=\mathrm{is\_similar}\left(o^{t}\rightarrow e^{t},o^{a}\right).$ (8)
Dual Reward.
Similar to the alignment reward, this reward also imposes constraints on $o^{t}$. Specifically, we require the emotion words extracted from $o^{t}$ to match the ground-truth label $y$:
$\mathcal{R}_{\text{dual}}(o|v,q,y)=\text{EW}\left(o^{t}\rightarrow e^{t},y\right).$ (9)
Perception Reward.
Aforementioned rewards primarily evaluate text-level outputs $o$ but overlook whether the generated output aligns with the input video $v$. To assess perception correctness, inspired by EmoPrefer [lian2025emoprefer], we employ MLLMs to estimate pairwise preferences among outputs $(o_{i},o_{j})$ for a given video. Then, we apply the Bradley-Terry algorithm [hunter2004mm] to derive the ranking from these preferences. Finally, we assign a reward of 1 to outputs ranked in the top 50% and 0 to the rest. This reward is defined as follows:
[…]
3.3.3 Reward Hacking
During reinforcement learning, we observe that the model tends to generate lengthy outputs $o^{a}$ containing numerous emotional words. The primary reason for this is that, in the calculation of EW-based metrics, synonyms are automatically disregarded (see Appendix A). Consequently, synonyms do not influence the metric scores, whereas predicting more words may instead increase the likelihood of matching the ground truth. This results in lengthy outputs, which in turn increases inference time and reduces the readability of $o^{a}$. This phenomenon, also referred to as reward hacking [everitt2017reinforcement], occurs when a model exploits loopholes or flaws in the reward function design to achieve high rewards while generating undesired behavior. To generate human-preferred outputs, we introduce length penalties based on the predicted label count, encouraging the model to maximize rewards with fewer prediction words. Specifically, we propose three types of penalties. Taking accuracy reward as an example, these penalties are defined as follows:
$P_{1}(o^{a},y)$ $=\begin{cases}1,&|o^{a}|\leq|y|\\ \frac{1}{|o^{a}|-|y|+1},&|o^{a}|>|y|\end{cases}$ (11)
$P_{2}(o^{a},y)$ $=\begin{cases}1,&|o^{a}|\leq|y|\\ \frac{|y|}{|o^{a}|},&|o^{a}|>|y|\end{cases}$ (12)
$P_{3}(o^{a},y)$ $=\begin{cases}1,&|o^{a}|\leq|y|\\ \frac{\log(|y|)}{\log(|o^{a}|)},&|o^{a}|>|y|\end{cases}$ (13)
where $|o^{a}|$ and $|y|$ denote the number of emotion words in the predictions and ground truth, respectively. We then apply these length penalties to the accuracy reward, yielding the final accuracy score:
$\mathcal{R}_{\text{accuracy}}(o|v,q,y)=P_{k}(o^{a},y)\times\text{EW}(o^{a},y).$ (14)
Similar to the accuracy reward, for the dual reward, we also apply this penalty:
$\mathcal{R}_{\text{dual}}(o|v,q,y)=P_{k}(o^{t}\rightarrow e^{t},y)\times\text{EW}(o^{t}\rightarrow e^{t},y).$ (15)
Figure 1 visualizes the values of these penalties (see the reward hacking part). From $P_{1}(\cdot)$ to $P_{3}(\cdot)$, the penalty on prediction lengths becomes looser. In Section 5.3, we conduct experiments to evaluate the effects of different penalty strategies on emotion recognition performance and their efficacy in mitigating redundant predictions.
4 Experimental Database and Setup
4.1 Corpus Description
AffectGPT-R1 contains two training phases: cold-start training and reinforcement learning. In the first phase, we employ a large-scale dataset to equip the model with emotion understanding and format alignment capabilities. For this purpose, we select MER-Caption+ [lian2025affectgpt], which contains 31K automatically annotated emotion descriptions. In the second phase, we need a fine-grained dataset to better address OV-MER. Thus, we choose MER2025-OV [lian2025mer], which includes 1,200 samples with human-annotated open-vocabulary emotion labels. For performance comparison, we evaluate the model on OV-MERD+ [lian2025affectgpt] and MER-UniBench [lian2025affectgpt], assessing its capabilities in open-vocabulary and generalized emotion understanding. Appendix D provides additional details about MER-UniBench. Since 200 samples overlap between MER2025-OV and OV-MERD+, we exclude these samples from MER2025-OV to ensure a fair comparison, resulting in 1,000 samples used in reinforcement learning. Table 1 presents statistics for these datasets. As shown, during cold-start training, we leverage both descriptive and open-vocabulary emotions provided by the dataset. However, in reinforcement learning, we exclusively use open-vocabulary emotions. This decision stems from the fact that the reward calculation process does not rely on human-annotated thinking results (see Section 3.3.2). Therefore, our dataset selection is well-suited to the distinct requirements of each training stage.
Table 1: Dataset statistics. We summarize the datasets used for training and testing. In this table, “Sen.”, “Bas.”, “OV”, and “Des.” are abbreviations for sentiment labels, basic emotions, open-vocabulary emotions, and descriptive emotions, respectively.
Dataset #Samples Task
Sen. Bas. OV Des.
Cold-start training
MER-Caption+ [lian2025affectgpt] 31,327 $\times$ $\times$ $\surd$ $\surd$
Reinforcement learning
MER2025-OV [lian2025mer] 1,000 $\times$ $\times$ $\surd$ $\times$
Testing
OV-MERD+ [lian2025affectgpt] 532 $\times$ $\times$ $\surd$ $\times$
MER-UniBench [lian2025affectgpt] 12,799 $\surd$ $\surd$ $\surd$ $\times$
4.2 Implementation Details
During training, we set the learning rate to 1e-5 and the maximum number of epochs to 60. The entire implementation is conducted using PyTorch, and the code is executed on an A100 GPU with 80 GB of memory. Due to memory constraints, we use a batch size of 3 for cold-start training and a batch size of 1 for reinforcement learning. For the model architecture, we adopt the default hyperparameters of AffectGPT [lian2025affectgpt]. This setup ensures a fair comparison between AffectGPT and AffectGPT-R1, better highlighting the effectiveness of our proposed framework.
5 Results and Discussion
In this section, we first analyze the effectiveness of different reward functions, the impact of the output mode, and the effect of varying length penalties. Next, we compare the performance of AffectGPT-R1 with other baselines on MER-UniBench, demonstrating the effectiveness of our method for generalized emotion understanding. Finally, we conduct a series of ablation studies to investigate the influence of the choice of $G$, as well as the amount of data used in cold-start training and reinforcement learning.
5.1 Reward Selection
This section evaluates the effectiveness of different reward functions and identifies the optimal reward strategy for addressing OV-MER. For comparison, we report performance under three scenarios: (1) without reinforcement learning, (2) with reinforcement learning using a single reward, and (3) with reinforcement learning using multiple rewards. Experimental results are presented in Table 2.
Single-reward RL.
In the single-reward scenarios, we observe that the accuracy reward and dual reward are the most effective, significantly improving performance compared to the model without reinforcement learning. In contrast, other rewards lead to performance degradation, indicating potential reward hacking. The superior performance of the accuracy reward and dual reward stems from their direct optimization of EW-based metrics: one on $o^{a}$ (answer) and one on $o^{t}$ (thinking). This highlights the importance of directly aligning reinforcement learning with evaluation metrics in OV-MER. Meanwhile, the format reward and alignment reward show limited impact, because these aspects are already well-learned during the cold-start phase. The weak performance of the perception reward may be due to current MLLMs’ inability to accurately decode human preferences, which could introduce reward noise [lian2025emoprefer]. Interestingly, the accuracy reward and dual reward yield similar performance, suggesting that while EW-based metrics are crucial, the choice of optimizing on thinking $o^{t}$ or answer $o^{a}$ has minimal impact. Therefore, in the next experiment, we use the accuracy reward by default.
Multiple-reward RL.
In the multiple-reward scenarios, we observe that combining the accuracy reward and format reward leads to further performance improvements compared to single-reward setups. However, using more rewards, such as all five simultaneously, results in performance degradation. The primary issue arises from interference between rewards: combining multiple objectives can dilute their individual effectiveness, ultimately harming reinforcement learning outcomes. For instance, emphasizing one reward may inadvertently reduce the influence of others. These findings suggest that more rewards do not necessarily equate to better performance. Instead, selecting an appropriate combination is key. Based on our experimental results, the most effective reward setup is the combination of the accuracy reward and format reward.
Table 2: Impact of reward function selection. In this table, we report performance under three conditions: (1) without reinforcement learning, (2) with reinforcement learning using a single reward, and (3) with reinforcement learning using multiple rewards.
Reward Functions OV-MERD+
accuracy format alignment dual preception
Without reinforcement learning
62.52
With reinforcement learning (single reward)
63.35(+0.83){(+0.83)}
59.22(3.30){(-3.30)}
57.55(4.97){(-4.97)}
64.98(+2.46){(+2.46)}
59.93(2.59){(-2.59)}
With reinforcement learning (multiple rewards)
66.49(+3.97){(+3.97)}
65.38(+2.86){(+2.86)}
64.66(+2.14){(+2.14)}
59.53(2.99){(-2.99)}
5.2 Output Mode Analysis
In single-reward setups (see Table 2), the most effective rewards are the accuracy reward and dual reward, both of which are directly optimized on EW-based metrics. This observation leads to a heuristic hypothesis: Can we modify the output mode to generate and optimize solely on $o^{a}$, thereby making the model more focused on EW-based metrics? Specifically, for cold-start training, we remove the thinking process and require the output to contain only $o^{a}$. For reinforcement learning, we optimize the accuracy reward on $o^{a}$ directly. We denote this modified model as AffectGPT-R1 (w/o think) and the original model in Figure 1 as AffectGPT-R1 (w/ think). Experimental results in Table 3 show that AffectGPT-R1 (w/o think) achieves better performance. This improvement can be attributed to two reasons. Firstly, our thinking process is derived from coarse-grained datasets, where dataset noise may introduce incorrect reasoning and degrade performance. Secondly, including additional thinking content diverts the model’s attention from addressing OV-MER, thereby increasing training difficulty.
Table 3: Output mode impact. This table reports the performance under different output types. We observe that a simpler model, one that only outputs answers and calculates rewards based solely on the answers, achieves better performance.
Model OV-MERD+
AffectGPT 62.52
AffectGPT-R1 (w/ think) 66.49
AffectGPT-R1 (w/o think) 68.39
5.3 Impact of Different Penalties
During inference, the model tends to generate lengthy responses with synonyms. This behavior primarily stems from the suboptimal design of the reward function. To mitigate this issue, we introduce three types of penalties in Section 3.3.3, where the penalty strength increases from $P_{3}(\cdot)$ to $P_{1}(\cdot)$. In this section, we evaluate the effectiveness of these penalties across three dimensions: (1) their impact on performance, (2) their ability to prevent overlong outputs, and (3) whether they introduce reward degradation. As shown in Table 4, all three penalties successfully address overlong outputs without compromising performance. However, the strict penalty, $P_{1}(\cdot)$, leads to reward degradation. Therefore, $P_{2}(\cdot)$ and $P_{3}(\cdot)$ are better choices.
Table 4: Impact of different penalties. The penalty strength increases from $P_{3}(\cdot)$ to $P_{1}(\cdot)$. Here, “overlong” indicates the presence of overlong outputs, and “degradation” indicates the presence of reward degradation. We seek a length penalty that preserves accuracy, prevents overlong outputs, and avoids causing reward degradation. Therefore, $P_{2}(\cdot)$ and $P_{3}(\cdot)$ are better choices.
Penalty Accuracy Overlong Degradation
$P_{1}(\cdot)$ 67.44
$P_{2}(\cdot)$ 67.79
$P_{3}(\cdot)$ 68.05
None 68.39
5.4 Main Results on MER-UniBench
The above sections primarily focus on open-vocabulary emotions. In this section, we broaden the evaluation scope to MER-UniBench [lian2025affectgpt], a unified benchmark encompassing three typical tasks: sentiment analysis, basic emotion recognition, and fine-grained emotion detection. Experimental results are presented in Table 5. For fairness, all comparisons are conducted under the same input modality. These results demonstrate that AffectGPT-R1 achieves notable performance improvements in basic and fine-grained emotion recognition but exhibits slight degradation in sentiment analysis. This discrepancy arises because fine-grained and basic MER primarily focus on discrete emotion words (e.g., happy), whereas sentiment analysis centers on emotion polarity (e.g., positive and negative). Consequently, fine-grained and basic MER are more closely related to each other. Based on the average results across three tasks, AffectGPT-R1 achieves state-of-the-art performance on MER-UniBench, demonstrating the advantages of reinforcement learning for multimodal emotion understanding.
Table 5: Main results on MER-UniBench. MER-UniBench is a unified benchmark comprising three key tasks: sentiment analysis, basic emotion recognition, and fine-grained emotion detection. We conduct a fair comparison among different baselines under identical input modalities. Overall, AffectGPT-R1 outperforms existing methods and achieves state-of-the-art performance in this benchmark.
Sentiment Basic Fine-grained Mean
MOSI MOSEI SIMS SIMS v2 MER2023 MER2024 MELD IEMOCAP OV-MERD+
Input Modality: Audio, Text
OneLLM [han2024onellm] 64.01 54.09 63.39 61.98 25.52 17.21 28.32 33.44 22.25 41.14
SECap [xu2024secap] 55.76 54.18 59.51 57.41 40.95 52.46 25.56 36.92 36.97 46.64
PandaGPT [su2023pandagpt] 66.06 61.33 62.93 58.88 33.57 39.04 31.91 36.55 31.33 46.84
Qwen-Audio [chu2023qwen] 70.09 46.90 70.73 65.26 41.85 31.61 49.09 35.47 32.36 49.26
SALMONN [tang2023salmonn] 81.00 67.03 68.69 65.93 55.53 45.38 45.62 46.84 45.00 57.89
AffectGPT [lian2025affectgpt] 83.46 80.74 82.99 83.75 72.94 73.41 56.63 55.68 59.98 72.18
AffectGPT-R1 80.13 80.01 84.49 82.31 81.69 93.49 63.74 63.85 65.49 77.24
Input Modality: Video, Text
Otter [li2025otter] 52.89 50.44 57.56 53.12 16.41 14.65 22.57 29.08 16.63 34.82
Video-LLaVA [lin2024video] 56.37 61.64 53.28 57.45 36.93 30.25 30.73 38.95 34.00 44.40
PandaGPT [su2023pandagpt] 58.50 64.25 62.07 65.25 39.13 47.16 38.33 47.21 35.07 50.77
Video-ChatGPT [maaz2024video] 54.42 63.12 64.82 65.80 44.86 46.80 37.33 56.83 39.80 52.64
VideoChat2 [li2024mvbench] 66.84 54.32 69.49 70.66 33.67 54.50 36.64 48.70 39.21 52.67
LLaMA-VID [li2024llama] 61.78 63.89 69.35 67.48 50.72 57.60 42.75 46.02 45.01 56.07
VideoChat [li2023videochat] 65.13 63.61 69.52 72.14 48.73 57.30 41.11 48.38 44.52 56.71
Chat-UniVi [jin2024chat] 54.53 63.18 68.15 66.36 57.62 65.67 45.61 52.37 48.00 57.94
mPLUG-Owl [ye2023mplug] 72.40 72.91 72.13 75.00 56.86 59.89 49.11 55.54 48.18 62.45
AffectGPT [lian2025affectgpt] 82.39 81.57 87.20 86.29 74.58 75.29 57.63 62.19 61.65 74.31
AffectGPT-R1 78.78 79.07 85.91 85.85 77.72 85.29 61.09 67.42 62.42 75.95
Input Modality: Audio, Video, Text
PandaGPT [su2023pandagpt] 61.92 67.61 68.38 67.23 40.21 51.89 37.88 44.04 37.12 52.92
R1-Omni [zhao2025r1] 58.02 56.48 71.82 68.58 64.17 67.43 43.20 51.58 55.24 59.61
Emotion-LLaMA [cheng2024emotion] 66.13 67.66 78.32 77.23 59.38 73.62 46.76 55.47 52.97 64.17
AffectGPT [lian2025affectgpt] 81.30 80.90 88.49 86.18 78.54 78.80 55.65 60.54 62.52 74.77
AffectGPT-R1 79.65 80.18 87.26 85.75 84.51 93.13 66.71 74.26 68.39 79.98
5.5 Ablation Study
Figure 2: Left: Impact of $G$. “–” denotes the baseline model without reinforcement learning. Mid: Impact of data amount in cold-start training. This figure presents the performance of models trained with varying amounts of cold-start data and analyzes their influence on subsequent reinforcement learning. Right: Impact of data amount in reinforcement learning. “–” denotes the baseline model without reinforcement learning. This figure shows that using a larger amount of data in reinforcement learning leads to better performance.
Impact of $G$.
GRPO requires sampling $G$ responses for group normalization (see Section 3.3.1). Figure 2 (left) reveals the impact of $G$. Due to GPU memory constraints, we limit our comparison to values of $G$ ranging from 2 to 16 and do not evaluate larger values. Experimental results demonstrate that, regardless of the chosen $G$, the model with reinforcement learning outperforms the one without reinforcement learning, demonstrating the effectiveness of the R1-style strategy in OV-MER. However, the selection of $G$ also matters. A small $G$ (e.g., $G=2$) fails to fully leverage the advantages of GRPO, primarily because insufficient sampling limits the likelihood of obtaining optimal results. Conversely, a larger $G$ (e.g., $G\geq 4$) ensures the effectiveness of GRPO, but further increasing $G$ (e.g., $G=16$) yields no significant improvement. Given that a larger $G$ also imposes higher GPU memory requirements, we adopt $G=4$ as our default setting.
Data Amount in Cold-start Training.
AffectGPT-R1 consists of two stages: cold-start training (denoted as $s_{1}$) and reinforcement learning (denoted as $s_{2}$). In this section, we investigate the impact of the amount of cold-start data on $s_{1}$ and $s_{2}$. Experimental results in Figure 2 (mid) demonstrate that increasing the amount of cold-start data improves performance in $s_{1}$. Meanwhile, we observe that the performance after reinforcement learning is closely tied to the performance of the base model from $s_{1}$, and more cold-start data also enhances $s_{2}$ performance. The main reason lies in that reinforcement learning predominantly serves to reinforce correct answers rather than infuse novel knowledge [liu2025understanding]. Therefore, a high-quality base model is essential for maximizing the effectiveness of reinforcement learning.
Data Amount in Reinforcement Learning.
This section investigates the influence of data quantity on reinforcement training. Specifically, we start from the same model obtained through cold-start training (i.e., the model trained on 100% data in Figure 2 (mid)) and evaluate its performance with different amounts of reinforcement learning data. Experimental results in Figure 2 (right) demonstrate that even using just 10 samples for reinforcement learning leads to performance improvements, and further increasing the data amount yields more substantial gains. These phenomenon indicates that AffectGPT-R1 has not yet reached its upper-bound performance. In the future, we plan to collect more data to further improve the performance of AffectGPT-R1.
6 Conclusion
This paper proposes AffectGPT-R1, the first work applying reinforcement learning to OV-MER. We introduce various rewards and incorporate length penalties to mitigate reward hacking. Through extensive experiments, we reveal the impact of different rewards and identify the optimal reward design. AffectGPT-R1 demonstrates strong effectiveness in OV-MER, and this framework also achieves state-of-the-art performance on MER-UniBench, a benchmark for generalized emotion understanding. Furthermore, we examine the influence of various modules and hyperparameters, providing valuable insights for future research. By proving the potential of reinforcement learning as a promising approach for emotion understanding, our work will advance the development of this field toward a reinforcement learning era.
[…]
Supplementary Material
A EW-based Metric
In this section, we present a detailed calculation formula for EW-based metrics. The computation method primarily consists of two components: eliminating the influence of synonyms and defining set-level metrics.
A.1 Handling Synonyms
To mitigate the impact of synonyms, we employ a three-level hierarchical grouping strategy:
(L1) We map different forms of words to their base form. For example, happier and happiness are mapped to happy. This function is denoted as $F_{l_{1}}(\cdot)$.
(L2) We map synonyms to a unified word. For example, happy and joyful are mapped to happy. This mapping function is denoted as $F_{l_{2}}(\cdot)$.
(L3) Emotion wheel provides natural grouping information [plutchik1980general]. Since there is no consensus on the emotion wheel, we adopt $K=5$ emotion wheels, consistent with prior work [lian2025ov]. Appendix B visualizes all emotion wheels. For each wheel $w_{k}$, we map all outer labels to their corresponding inner labels. This function is denoted as $F_{l_{3}}^{w_{k}}(\cdot)$.
The clustering functions can be summarized as:
$G_{w_{k}}(\cdot)=F_{l_{3}}^{w_{k}}{\left(F_{l_{2}}\left(F_{l_{1}}\left(\cdot\right)\right)\right)},k\in[1,K].$ (16)
A.2 Set-level Metric
Since we do not restrict the number of predicted and annotated emotions, we employ set-level evaluation metrics. Specifically, suppose the dataset contains $N$ samples. For a sample $x_{i}$, the true labels are $\mathbf{Y}_{i}=\{y_{i}^{j}\}_{j=1}^{n_{i}}$ and the predicted labels are $\mathbf{\hat{Y}}_{i}=\{\hat{y}_{i}^{j}\}_{j=1}^{\hat{n}_{i}}$. The evaluation metrics are defined as follows:
$\mbox{Precision}_{\mbox{s}}^{k}=\frac{1}{N}\sum_{i=1}^{N}\frac{\left|G_{w_{k}}(\mathbf{Y}_{i})\cap G_{w_{k}}(\mathbf{\hat{Y}}_{i})\right|}{\left|G_{w_{k}}(\mathbf{\hat{Y}}_{i})\right|},$ (17)
$\mbox{Recall}_{\mbox{s}}^{k}=\frac{1}{N}\sum_{i=1}^{N}\frac{\left|G_{w_{k}}(\mathbf{Y}_{i})\cap G_{w_{k}}(\mathbf{\hat{Y}}_{i})\right|}{\left|G_{w_{k}}(\mathbf{{Y}}_{i})\right|},$ (18)
$\mbox{F}_{\mbox{s}}^{k}=2\times\frac{\mbox{Precision}_{\mbox{s}}^{k}\times\mbox{Recall}_{\mbox{s}}^{k}}{\mbox{Precision}_{\mbox{s}}^{k}+\mbox{Recall}_{\mbox{s}}^{k}}.$ (19)
In this set-wise operation, we automatically remove duplicate emotion words. Finally, we compute the average F1-score across all emotion wheels as the final score:
$\text{EW}\left(\mathbf{Y}_{i},\mathbf{\hat{Y}}_{i}\right)=\frac{1}{K}\sum_{k=1}^{K}\mbox{F}_{\mbox{s}}^{k}.$ (20)
B Emotion Wheel
The emotion wheel is a structured visual tool designed to categorize and identify human emotions. It organizes emotions into core categories (the inner sections of the wheel) and their nuanced variations (the outer sections), providing intuitive insights into emotional structure. This paper uses five emotion wheels to compute metrics for OV-MER, aligning with prior work [lian2025affectgpt] to ensure fair comparisons.
(a) W1
(b) W2
(c) W3
(d) W4
(e) W5
Figure 3: Emotion wheels. This paper uses five emotion wheels to calculate EW-based metrics.
C Baseline Details
SECap [xu2024secap].
This is a speech emotion captioning framework designed to describe speech emotions using natural language. Specifically, it employs HuBERT as the audio encoder to extract speech representations and then utilizes the Bridge-Net to obtain emotion-relevant speech features. Finally, these features are fed into the LLM, which generates natural language captions describing speech emotions.
SALMONN [tang2023salmonn].
This model is capable of processing both speech and audio. Specifically, it employs Whisper [radford2023robust] to encode speech and BEATs [chen2023beats] to encode audio. The resulting features are then integrated into the LLM via a window-level Q-Former. Additionally, a LoRA adapter is applied to enhance the LLM’s instruction-following and question-answering capabilities.
Qwen-Audio [chu2023qwen].
This model is capable of handling 30 different tasks and a variety of audio types, demonstrating universal audio understanding capabilities. Specifically, it employs an audio encoder to map input audio into hidden features. To address interference between different tasks, it adopts a multi-task input format. The output features are then passed to LLMs for answer generation.
Otter [li2025otter].
This model is built upon the OpenFlamingo framework [awadalla2023openflamingo] and trained on the MIMIC-IT dataset [li2025otter], demonstrating strong capabilities in multimodal perception, reasoning, and in-context learning.
Chat-UniVi [jin2024chat].
This is a unified vision-language model capable of processing both images and videos. It employs dynamic visual tokens that capture spatial details in images and temporal relationships in videos.
Video-LLaVA [lin2024video].
It aligns images and videos before projection, enabling the LLM to learn from a unified visual representation and equipping it with the ability to comprehend both images and videos simultaneously.
Video-ChatGPT [maaz2024video].
This model averages frame-level features across temporal and spatial dimensions, extracting both spatial and temporal video features. These features are then projected into the text token space, enabling the LLM to generate responses.
VideoChat [li2023videochat].
It uses video foundation models to encode videos as embeddings, then employs LLMs for response generation. After instruction fine-tuning, the model demonstrates promising capabilities across various video applications.
VideoChat2 [li2024mvbench].
This model leverages diverse instruction-tuning data and employs a three-stage progressive multimodal training approach, comprising vision-language alignment, vision-language connection, and instruction tuning. As a result, it outperforms leading models across multiple tasks.
LLaMA-VID [li2024llama].
This model represents each image with two tokens: a context token and a content token. The context token encodes the global image information based on user input, while the content token captures visual details within the image. By condensing each image into just a few tokens, the model can efficiently understand and process long videos.
mPLUG-Owl [ye2023mplug].
This approach equips LLMs with vision-language capabilities through a two-stage training paradigm, enabling strong performance on instruction-following, visual understanding, and reasoning tasks.
PandaGPT [su2023pandagpt].
It uses ImageBind [girdhar2023imagebind] to map multiple modalities into a shared embedding space, then employs a projector to transform these features into the LLM’s input space, and finally leverages the LLM to generate responses.
OneLLM [han2024onellm].
This model aligns eight modalities with language through a unified framework. Its architecture consists of a universal encoder, a universal projection module, and modality-specific tokens. Additionally, it employs a progressive multimodal alignment pipeline, enabling it to comprehend eight input types and generate responses based on user queries.
Emotion-LLaMA [cheng2024emotion].
This framework follows the architecture of traditional MLLMs and consists of three key components: encoders, projectors, and LLMs. Specifically, it inputs text, audio, peak frames, and three levels of visual cues (i.e., local, temporal, and global representations) and employs modality-specific encoders to convert these inputs into corresponding hidden features. These features are then aligned into a shared dimensional space with text embeddings through a linear projection mechanism. Finally, the LLM integrates information from all modalities for emotion recognition and reasoning. Experimental results on benchmark datasets demonstrate that this framework achieves improved performance over previous solutions.
R1-Omni [zhao2025r1].
It focuses on basic emotion recognition and uses reinforcement learning with verifiable rewards to address this task. Unlike R1-Omni, AffectGPT-R1 shifts the task to open-vocabulary emotion detection. Due to this task difference, we adopt distinct reward functions and face unique challenges, such as reward hacking, that are not present in R1-Omni. It is important to note that R1-Omni is only available as a preprint on arXiv. Following CVPR policy, we are not required to compare with it. Nevertheless, to thoroughly validate the effectiveness of AffectGPT-R1, we include a comparison with R1-Omni, and the experimental results demonstrate our model’s advantages in open-vocabulary emotion detection and generalized emotion understanding. Additionally, unlike R1-Omni, which only releases its inference code, we provide both training and inference code to enhance reproducibility.
AffectGPT [lian2025affectgpt].
This framework differs from previous MLLMs, which typically delegate the entire multimodal fusion process to the language model. Instead, it relocates cross-modal interaction outside the language model and incorporates a pre-fusion operation to enhance multimodal integration, emphasizing the inherently multimodal nature of human emotions. Specifically, the framework introduces two types of pre-fusion mechanisms: Q-Former and attention mechanisms. Both modules are computationally efficient and introduce only a small number of trainable parameters. Experimental results demonstrate the effectiveness of AffectGPT in generalized emotion understanding, including sentiment analysis, basic emotion recognition, and fine-grained emotion detection.
D MER-UniBench Details
MER-UniBench is a unified benchmark that covers three typical tasks, including sentiment analysis, basic emotion recognition, and fine-grained emotion detection, with the aim of evaluating the generalized emotion understanding capabilities of different models. In Table 6, we summarize the data components for each task. Please refer to the original paper for more details [lian2025affectgpt].
Table 6: MER-UniBench. This benchmark encompasses three tasks, and this table presents the data components for each task. In total, this benchmark includes 12,799 samples for testing.
Raw Dataset Selected Subset # Samples
Sentiment Analysis
CMU-MOSI [zadeh2017tensor] Test 686
CMU-MOSEI [zadeh2018multimodal] Test 4,659
CH-SIMS [yu2020ch] Test 457
CH-SIMS v2 [liu2022make] Test 1,034
Basic Emotion Recognition
MER2023 [lian2023mer] MER-MULTI 411
MER2024 [lian2024mer] MER-SEMI 1,169
IEMOCAP [busso2008iemocap] Session5 1,241
MELD [poria2019meld] Test 2,610
Fine-grained Emotion Detection
OV-MERD+ [lian2025affectgpt] All 532
E Impact Statement
E.1 Social Impact
Emotion plays an important role in mental health, human-computer interaction, and embodied AI. This paper aims to advance the development of MER techniques, thereby promoting their applications and improving our daily lives.
E.2 Ethics Statement
This paper does not collect or annotate new data but utilizes existing datasets to verify the effectiveness of reinforcement learning in OV-MER. The use of all datasets has been authorized by their owners under corresponding usage requirements. Therefore, no ethical issues are raised in this paper.
F Reproducibility Statement
In this paper, we have made every effort to ensure the reproducibility of our work. In the supplementary material, we provide our source code and command lines for cold-start training, reinforcement learning, and inference. Additionally, we include a detailed README file. To comply with CVPR’s anonymity requirements, we have removed all external links and owner names. After acceptance, we will release the full code on GitHub to facilitate further research.

### v3

@@
 Abstract
-Open-Vocabulary Multimodal Emotion Recognition (OV-MER) aims to predict emotions without being constrained by predefined label spaces, enabling fine-grained emotion understanding.
-Unlike traditional discriminative methods, OV-MER leverages generative models, such as large language models, to capture the full spectrum of emotions and employs emotion wheels (EWs) for metric calculation.
+Open-Vocabulary Multimodal Emotion Recognition (OV-MER) aims to predict emotions without being constrained by label spaces, enabling fine-grained emotion understanding.
+Unlike traditional discriminative methods, OV-MER leverages generative models to capture the full spectrum of emotions and employs emotion wheels (EWs) for metric calculation.
 Previous approaches (e.g., AffectGPT) primarily rely on token-level loss during training.
 However, this objective is misaligned with the metrics used in OV-MER, while these metrics cannot be optimized via gradient backpropagation.
-In this paper, we propose AffectGPT-R1, a reinforcement learning framework that formulates EW-based metrics as a reward function and employs a policy-based optimization strategy to maximize this reward.
-Additionally, we introduce an extra reasoning process and investigate its necessity in OV-MER.
-To further refine model behavior, we incorporate auxiliary rewards that constrain both reasoning and emotion prediction.
-To prevent reward hacking, we propose to incorporate length penalties during training.
-Experimental results show that AffectGPT-R1 achieves substantial improvements on OV-MER.
-Beyond this task, our approach also enhances generalized emotion understanding, attaining state-of-the-art performance on MER-UniBench.
-To the best of our knowledge, this is the first work to adapt the R1-style methodology for emotion understanding, revealing the impact of reasoning processes and reinforcement learning in this domain.
+To address this limitation, we propose AffectGPT-R1, a reinforcement learning framework that treats EW-based metrics as a reward function and applies policy optimization to maximize this reward.
+Additionally, we introduce an explicit reasoning process and examine its necessity in OV-MER.
+To further guide model behavior, we incorporate auxiliary rewards that regularize both emotion reasoning and emotion prediction.
+We also apply length penalties to mitigate reward hacking.
+Experimental results demonstrate that AffectGPT-R1 yields significant performance improvements on OV-MER.
+Moreover, our approach enhances generalized emotion understanding, achieving state-of-the-art results on MER-UniBench.
 Our code is provided in the supplementary material and will be released to facilitate future research.
+Machine Learning, ICML
 1 Introduction
-Emotion is highly related to human cognition, decision-making, and behavior, playing a vital role in our daily life [picard2000affective, xie2024emovit].
-Naturally, humans express emotions through multiple modalities, including (micro)-gestures, (micro)-expressions, audio tones, linguistic content, and other cues [ben2021video, chen2023smg, el2011survey, lian2024merbench].
+Emotion is highly related to human cognition, decision-making, and behavior, playing a vital role in our daily life (Picard, 2000; Xie et al., 2024).
+Naturally, humans express emotions through multiple modalities, including (micro)-gestures, (micro)-expressions, audio tones, linguistic content, and other cues (Ben et al., 2021; Chen et al., 2023a; El Ayadi et al., 2011; Lian et al., 2026b).
 To better understand human emotions, it is essential to integrate these clues, enabling models to capture emotional nuances.
-This has led to the development of Multimodal Emotion Recognition (MER) [lian2023mer, zhang2025moda].
-Traditional MER approaches primarily rely on Ekman’s theory [ekman1992argument, li2023decoupled], which categorizes human emotions into six basic labels: anger, disgust, happiness, sadness, fear, and surprise.
-However, human emotions extend far beyond these basic categories, and such a limited label space inevitably results in inaccurate emotion descriptions [plutchik1980general, demszky2020goemotions].
-Recently, open-vocabulary MER (OV-MER) has emerged as a promising research direction, aiming to shift emotion recognition from constrained categories to the full spectrum of emotion, thereby enabling more nuanced emotion representations [lian2025ov, lian2025affectgpt].
+This has led to the development of Multimodal Emotion Recognition (MER) (Lian et al., 2023a; Zhang et al., 2025b).
+Traditional MER approaches primarily rely on Ekman’s theory (Ekman, 1992; Li et al., 2023), which categorizes human emotions into six basic labels: anger, disgust, happiness, sadness, fear, and surprise.
+However, human emotions extend far beyond these basic categories (Plutchik, 1980; Demszky et al., 2020), and such a restricted label space inevitably leads to imprecise emotion descriptions, thereby impairing the emotional intelligence of developed systems.
+Recently, open-vocabulary MER (OV-MER) has emerged as a promising research direction, aiming to shift emotion recognition from constrained categories to the full spectrum of emotion, thereby enabling more nuanced emotion representations (Lian et al., 2025a, c).
 To support this paradigm shift, OV-MER introduces new solutions and metrics.
 For solutions, it transitions from discriminative to generative approaches, leveraging the extensive vocabulary of large language models (LLMs) to expand the recognition scope of emotion labels.
-For metrics, it employs emotion wheels (EWs) to capture semantic relationships between distinct emotion words.
-Appendix A provides a detailed illustration of these metrics.
-To better address this task, current approaches primarily use token-level loss to align predicted and ground-truth labels [lian2025affectgpt, cheng2024emotion].
+For metrics, it employs emotion wheels (EWs) to capture semantic relationships between distinct emotions.
+Appendix D provides a detailed illustration of these metrics.
+To better address this task, current approaches primarily use token-level loss to align predicted and true labels (Lian et al., 2025a; Cheng et al., 2024).
 However, these methods suffer from a critical misalignment: token-level loss shows limited correlation with EW-based metrics.
@@
 To refine model behaviors, we incorporate auxiliary reward functions to constrain the model’s reasoning and emotion predictions.
-Through extensive experiments, we analyze the impact of different reward functions and identify the most effective ones.
-Additionally, during inference, the reward hacking problem [everitt2017reinforcement] arises, where the model tends to generate lengthy predictions with redundant emotional expressions (e.g., happy and happily).
-To mitigate this, we incorporate length penalties into the rewards, ensuring reward maximization while promoting more concise outputs.
-Experimental results demonstrate that AffectGPT-R1 achieves significant performance improvements on OV-MER.
-Our framework also attains state-of-the-art performance on MER-UniBench [lian2025affectgpt], a unified benchmark for generalized emotion understanding, further validating its effectiveness.
-To the best of our knowledge, this is the first work to apply R1-style training to OV-MER, demonstrating the effect of reasoning processes and reinforcement learning in this domain.
-The main contributions of this paper are summarized as follows:
-  • (Framework) We introduce AffectGPT-R1, the first work to apply an R1-style training strategy to OV-MER, shedding light on the role of reinforcement learning and reasoning processes in this task.
-  • (Reward) We systematically investigate various reward functions and conduct an in-depth analysis of their effectiveness, identifying the optimal rewards for OV-MER.
-To mitigate the reward hacking problem, we introduce length penalties, ensuring the model maximizes rewards while using the fewest emotional words.
-  • (Performance) We achieve state-of-the-art performance on OV-MER and MER-UniBench, proving the potential of reinforcement learning for emotion understanding.
+Additionally, we observe the issue of reward hacking (Everitt et al., 2017), wherein the model tends to produce lengthy predictions containing redundant emotional expressions.
+To address this, we introduce length penalties into the reward function, thereby encouraging more concise outputs while still maximizing overall reward.
+Experimental results demonstrate that AffectGPT-R1 achieves significant improvements on OV-MER.
+Our framework also attains state-of-the-art performance on MER-UniBench (Lian et al., 2025a), a unified benchmark for generalized emotion understanding.
+Appendix B provides a more detailed discussion of our novelty.
+The main contributions of this paper are summarized below:
+  • (Framework) We introduce AffectGPT-R1, a pioneering work that applies reinforcement learning to OV-MER, shedding light on the role of reinforcement learning and reasoning mechanisms in this domain.
+  • (Reward) We systematically investigate various reward functions and conduct an in-depth analysis of their effectiveness.
+To mitigate reward hacking, we introduce length penalties, ensuring the model maximizes rewards while using the fewest emotional words.
+  • (Performance) We achieve state-of-the-art performance in open-vocabulary emotion detection and generalized emotion understanding, highlighting the potential of reinforcement learning in affective computing.
 2 Related Works
 2.1 Emotion Representation
-Emotion representation defines the recognition scope of developed models and serves as the foundation of MER [gunes2011emotion].
-There are two primary representation methods: dimensional emotion theory [schlosberg1954three] and basic emotion theory [ekman1992argument].
+Emotion representation defines the recognition scope of developed models and serves as the foundation of MER (Gunes et al., 2011).
+There are two primary representation methods: dimensional emotion theory (Schlosberg, 1954) and basic emotion theory (Ekman, 1992).
 Dimensional Emotion Theory.
 This theory uses continuous dimensions to capture the core attributes underlying emotions.
-For example, Russell [russell1980circumplex] proposed a two-dimensional model, characterizing emotions based on valence (the pleasantness of an emotion) and arousal (the activation level of an emotion).
-In addition to this, some researchers introduced an additional dominance dimension to reflect the perceived control over emotions [russell1977evidence].
-However, these dimensions are abstract and less intuitive for the general public, posing challenges for applications.
+For example, Russell (Russell, 1980) proposed a two-dimensional model, characterizing emotions based on valence (the pleasantness of an emotion) and arousal (the activation level of an emotion).
+In addition, some researchers have introduced a third dimension, dominance, to represent the degree of perceived control over emotions (Russell and Mehrabian, 1977).
+However, these dimensions are abstract and less intuitive for the general public, thus posing challenges in practical applications.
 Basic Emotion Theory.
-This theory typically uses six or eight basic labels to describe emotional states.
-For example, Ekman [ekman1992argument] argued that human emotions can be categorized into six basic labels, while Plutchik [plutchik1980general] proposed eight primary emotions that serve as the foundation for all others.
+This theory typically relies on six or eight basic labels to describe emotional states.
+For example, Ekman (Ekman, 1992) argued that human emotions can be categorized into six basic labels, whereas Plutchik (Plutchik, 1980) identified eight primary emotions that form the basis for other emotional experiences.
 Nevertheless, human emotions are rich and diverse, extending far beyond these basic labels.
@@
 Open-vocabulary Emotion.
-To address these limitations, researchers propose an open-vocabulary representation manner [lian2025ov], which removes restrictions on the label space and enables models to predict any emotion.
+To address these limitations, researchers propose an open-vocabulary representation manner (Lian et al., 2025c), which removes restrictions on the label space and enables models to predict any emotion.
 Unlike dimensional emotion theory, it focuses on more intuitive emotion words; Unlike basic emotion theory, it avoids restrictive label spaces, enabling more nuanced emotion representation.
-This paper focuses on open-vocabulary emotions and introduces AffectGPT-R1, a pioneering work that leverages the R1-style training strategy to tackle this task.
+Therefore, the open-vocabulary paradigm opens up new opportunities for affective computing.
+However, this field is still in its early stages, and effective solutions remain unclear.
+This paper presents a pioneering effort that leverages reinforcement learning to tackle this task.
 Figure 1: Overall pipeline of AffectGPT-R1.
@@
 In the first phase, we adopt the framework of AffectGPT but replace the generated content, including both thinking and answers.
-In the second phase, we propose five rewards and introduce length penalties to mitigate reward hacking, followed by GRPO for policy updates.
+In the second phase, we propose five rewards and introduce length penalties to mitigate reward hacking.
 2.2 Reinforcement Learning
-Reinforcement learning has advanced the upper-bound performance of LLMs, demonstrating impressive results in tasks such as mathematical verification and code execution [guo2025deepseek, shao2024deepseekmath, cobbe2021training].
+Reinforcement learning has advanced the upper-bound performance of LLMs, demonstrating impressive results in tasks such as mathematical verification and code execution (Guo et al., 2025; Shao et al., 2024; Cobbe et al., 2021).
 This method primarily consists of two key components: policy optimization and reward modeling.
@@
 During training, we aim to maximize the reward function through policy updates.
-Proximal Policy Optimization (PPO) [schulman2017proximal] is a widely used algorithm that leverages a value model for advantage estimation and incorporates token-wise KL penalties to prevent excessive optimization.
-Direct Preference Optimization (DPO) [rafailov2023direct] focuses on directly optimizing policies to align with human preferences, eliminating the need for an explicit reward model (as required in PPO) and instead relying on an implicit reward model.
+Proximal Policy Optimization (PPO) (Schulman et al., 2017) is a widely used algorithm that leverages a value model for advantage estimation and incorporates token-wise KL penalties to prevent excessive optimization.
+Direct Preference Optimization (DPO) (Rafailov et al., 2023) focuses on directly optimizing policies to align with human preferences, eliminating the need for an explicit reward model and instead relying on an implicit reward model.
 This approach reduces computational costs and engineering complexity.
-Group Relative Policy Optimization (GRPO) [shao2024deepseekmath] is also a more efficient alternative to PPO, removing the value model and instead using a group-relative baseline computed from different outputs for the same input to guide optimization.
-Beyond these algorithms, some efforts have been made to enhance training efficiency and stability [hu2025reinforce++, yu2025dapo].
-In this paper, we do not focus on designing novel policy optimization algorithms.
-Instead, we adopt GRPO for policy updates.
+Group Relative Policy Optimization (GRPO) (Shao et al., 2024) is a more efficient alternative to PPO, removing the value model and instead using a group-relative baseline computed from different outputs for the same input to guide optimization.
+Beyond these algorithms, some efforts have been made to enhance training efficiency and stability (Hu, 2025; Yu et al., 2025).
+This paper does not focus on designing new policy optimization algorithms, but rather on crafting optimal rewards for emotion recognition.
 Reward Modeling.
@@
 The former is suitable for tasks with verifiable outcomes.
-For example, pass or fail unit tests in code generation [le2022coderl], output language requirements in question-answering tasks [guo2025deepseek], and IoU between predicted bounding boxes and ground-truth annotations in object detection [shen2025vlm].
-The latter is mainly used for tasks where rewards are difficult to compute, such as safety, helpfulness, and harmlessness [zhang2024safetybench].
+For example, pass or fail unit tests in code generation (Le et al., 2022), output language requirements in question-answering tasks (Guo et al., 2025), and IoU between predicted bounding boxes and ground-truth annotations in object detection (Shen et al., 2025).
+The latter is mainly used for tasks where rewards are difficult to compute, such as safety, helpfulness, and harmlessness (Zhang et al., 2024).
 In prior work, how to design reward functions tailored for OV-MER remains unexplored.
-To this end, we introduce multiple rewards to supervise both thinking and answers, deriving insights into the effective rewards for OV-MER.
+To this end, we introduce multiple rewards to supervise both thinking and answers, deriving insights into the effective rewards for this task.
 3 AffectGPT-R1
 Our training pipeline consists of two phases: cold start training and reinforcement learning.
-In the first phase, we leverage a coarse-grained dataset to establish initial capabilities in emotion understanding and format alignment.
-In the second phase, we use a fine-grained dataset for reward calculation and policy updates.
-Figure 1 illustrates the overall training pipeline of AffectGPT-R1.
+In the first phase, we utilize a large-scale, coarse-grained dataset to establish initial capabilities in emotion understanding and format alignment.
+In the second phase, we use a high-quality, fine-grained dataset for reward calculation and policy updates.
+Figure 1 illustrates the overall pipeline of AffectGPT-R1.
 3.1 Notation Definition
@@
 Figure 1 clarifies the meanings of these notations.
-In this paper, our main goal is to predict $y$ based on the video content $v$.
+Our main goal is to predict the emotion $y$ based on the video content $v$.
+We would like to emphasize that this paper focuses on utterance-level emotion recognition, ensuring fair comparison with prior work (Lian et al., 2025a, c).
+Appendix G.4 provides additional clarification regarding our task formulation.
 3.2 Cold Start Training
-Human emotions are highly dependent on subtle cues [li2022deep, li2022cas], posing significant challenges for current models to capture all these signals and accurately interpret emotional states [lian2024gpt, lu2024gpt].
+Human emotions are highly dependent on subtle cues (Li et al., 2022a, b), posing significant challenges for current models to capture all these signals and accurately interpret emotional states (Lian et al., 2024b; Lu et al., 2024).
 To enhance emotion understanding, we employ large-scale coarse-grained descriptive emotion datasets for cold-start training.
-In this phase, we adopt the same model architecture as AffectGPT [lian2025affectgpt] (see Figure 1).
+In this phase, we adopt the same model architecture as AffectGPT (Lian et al., 2025a) (see Figure 1).
 This framework leverages a pre-fusion layer to effectively integrate multimodal cues, emphasizing the inherently multimodal nature of human emotions.
 Unlike the original AffectGPT, which only outputs emotion words, we generate both thinking (enclosed within <<think>> and <</think>> tags) and answers (enclosed within <<answer>> and <</answer>> tags).
-Since descriptive emotion datasets already follow a thinking-like format [lian2023explainable], we treat these descriptions as the thinking content.
+Since descriptive emotion datasets already follow a thinking-like format (Lian et al., 2023b), we treat these descriptions as the thinking content.
 Given our primary focus on OV-MER, we use open-vocabulary emotions as the answer content.
@@
 3.3.1 Policy Optimization
-We employ GRPO [shao2024deepseekmath] for policy updates.
-Unlike PPO [schulman2017proximal], which relies on a critic model for advantage estimation, GRPO eliminates this component and instead uses group scores, significantly reducing memory overhead while improving training efficiency.
-Specifically, for a given input pair $(v,q)$, the policy model $\pi_{\theta_{\text{old}}}$ samples a group of responses $\mathbf{O}=\{o_{1},\cdots,o_{G}\}$, where $o_{i}$ is the $i$-th response.
+We employ GRPO (Shao et al., 2024) for policy updates.
+Unlike PPO (Schulman et al., 2017), which relies on a critic model for advantage estimation, GRPO eliminates this component and instead uses group scores, significantly reducing memory overhead while improving training efficiency.
+Specifically, for a given input pair $(v,q)$, the policy model $\pi_{\theta_{\text{old}}}$ samples a group of responses $\mathbf{O}=\{o_{i}\}_{i=1}^{G}$.
 We then compute the reward for each response as $r_{i}=\mathcal{R}(o_{i}|v,q,y)$, where $\mathcal{R}(\cdot)$ is the reward function defined in Section 3.3.2.
-To evaluate the relative quality of $\{o_{i}\}_{i=1}^{G}$, we normalize the rewards by computing their mean and standard deviation, deriving the advantage as:
+To evaluate their relative quality, we normalize the rewards by computing their mean and standard deviation:
 $\hat{A}_{i}=\frac{r_{i}-\text{mean}(\{r_{i}\}_{i=1}^{G})}{\text{std}(\{r_{i}\}_{i=1}^{G})}.$ (1)
@@
 The combination of clipping and KL regularization effectively prevents extreme policy updates while maintaining training stability.
-During experiments, we observed significant divergence between the learned policy distribution and the initial model distribution, and the restrictive effect of the KL penalty term led to performance degradation.
-This trade-off between policy exploration and regularization has also been proven in prior work [everitt2017reinforcement].
-To address this limitation, we remove the KL regularization term from our algorithm and adopt the following simplified objective function for policy update:
+During experiments, we observed a divergence between the learned policy distribution and the initial model distribution, and the restrictive nature of the KL penalty term constrained overall performance.
+This trade-off between policy exploration and regularization has been proven in prior work (Yu et al., 2025).
+Following the prior work (Yu et al., 2025), we remove the KL regularization term and adopt the simplified objective function for policy update.
+Experimental results in Section 5 demonstrate the effectiveness of our method.
 $\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}_{(v,q)\sim\mathcal{D},\{o_{i}\}_{i=1}^{G}\sim\pi_{\theta_{\text{old}}}(\cdot|v,q)}\frac{1}{G}\sum_{i=1}^{G}S_{i}.$ (5)
 3.3.2 Reward Modeling
-Reward functions guide the update direction of policy models.
-In this work, we propose five reward functions, all designed to evaluate thinking and answers.
-In Section 5.1, we examine the performance of the single reward and combinations of multiple rewards.
-Our goal is to identify the most effective reward strategy for OV-MER.
+Reward functions guide the update direction.
+This paper proposes five reward functions, all designed to steer thinking and answers.
+Section 5.1 reveals the performance of the single reward and combinations of multiple rewards.
 Format Reward.
@@
 Accuracy Reward.
-We employ the official metrics of OV-MER to calculate accuracy rewards [lian2025ov].
-Unlike discriminative MER, which operates within a fixed label space, OV-MER does not constrain the label space.
-Therefore, models are permitted to generate synonyms, i.e., emotions with the same meaning but expressed differently, which introduces challenges in evaluation.
-To mitigate the impact of synonyms, OV-MER introduces EW-based metrics, leveraging the structural information of emotion wheels for clustering.
-In these wheels, the inner section represents core emotions, while the outer section captures more nuanced emotions.
-Calculation details can be found in Appendix A:
+We employ the official metrics of OV-MER to compute accuracy rewards (Lian et al., 2025c).
+Unlike discriminative MER, which operates within a fixed label space, OV-MER imposes no constraints on the label space.
+Consequently, models are allowed to generate synonyms, i.e., emotions with identical meanings but different expressions, which poses challenges for evaluation.
+To mitigate these challenges, OV-MER introduces EW-based metrics, which leverage the structural information of emotion wheels for clustering.
+Following prior work (Lian et al., 2025c), we utilize five emotion wheels (see Appendix C), and a detailed calculation procedure for EW-based metrics is presented in Appendix D.
+We denote this reward as:
 $\mathcal{R}_{\text{accuracy}}(o|v,q,y)=\text{EW}\left(o^{a},y\right).$ (7)
 Alignment Reward.
-The accuracy reward ensures the correctness of $o^{a}$.
-However, without constraints on the thinking process, the generated output may contain unrelated $o^{a}$ and $o^{t}$.
-To this end, we introduce the alignment reward, ensuring that the reasoning process aligns with the final answer.
-Specifically, for each output, we extract emotion words $e^{t}$ from $o^{t}$ and compute their similarity with $o^{a}$:
+Accuracy reward ensures the correctness of $o^{a}$.
+However, without constraints on the reasoning process, the generated output may include irrelevant $o^{a}$ and $o^{t}$.
+To address this, we introduce alignment reward, which ensures that the reasoning process is consistent with the final answer.
+Specifically, we extract emotion words $e^{t}$ from $o^{t}$ and compute their similarity with $o^{a}$.
+Appendix E provides the detailed methods for emotion word extraction and similarity calculation.
+We denote this reward as:
 $\mathcal{R}_{\text{alignment}}(o|v,q,y)=\mathrm{is\_similar}\left(o^{t}\rightarrow e^{t},o^{a}\right).$ (8)
 Dual Reward.
-Similar to the alignment reward, this reward also imposes constraints on $o^{t}$. Specifically, we require the emotion words extracted from $o^{t}$ to match the ground-truth label $y$:
+Similar to alignment reward, this reward also imposes constraints on $o^{t}$. Specifically, we require the emotion words extracted from $o^{t}$ to match the true label $y$:
 $\mathcal{R}_{\text{dual}}(o|v,q,y)=\text{EW}\left(o^{t}\rightarrow e^{t},y\right).$ (9)
@@
 Aforementioned rewards primarily evaluate text-level outputs $o$ but overlook whether the generated output aligns with the input video $v$.
-To assess perception correctness, inspired by EmoPrefer [lian2025emoprefer], we employ MLLMs to estimate pairwise preferences among outputs $(o_{i},o_{j})$ for a given video.
-Then, we apply the Bradley-Terry algorithm [hunter2004mm] to derive the ranking from these preferences.
+To assess perception correctness, inspired by EmoPrefer (Lian et al., 2026a), we employ MLLMs to estimate pairwise preferences among outputs $(o_{i},o_{j})$ for a given video.
+Then, we apply the Bradley-Terry algorithm (Hunter, 2004) to derive the ranking from these preferences.
 Finally, we assign a reward of 1 to outputs ranked in the top 50% and 0 to the rest.
@@
 During reinforcement learning, we observe that the model tends to generate lengthy outputs $o^{a}$ containing numerous emotional words.
-The primary reason for this is that, in the calculation of EW-based metrics, synonyms are automatically disregarded (see Appendix A).
+The primary reason for this is that, in the calculation of EW-based metrics, synonyms are automatically disregarded (see Appendix D).
 Consequently, synonyms do not influence the metric scores, whereas predicting more words may instead increase the likelihood of matching the ground truth.
 This results in lengthy outputs, which in turn increases inference time and reduces the readability of $o^{a}$.
-This phenomenon, also referred to as reward hacking [everitt2017reinforcement], occurs when a model exploits loopholes or flaws in the reward function design to achieve high rewards while generating undesired behavior.
+This phenomenon, also referred to as reward hacking (Everitt et al., 2017), occurs when a model exploits loopholes or flaws in the reward function design to achieve high rewards while generating undesired behavior.
 To generate human-preferred outputs, we introduce length penalties based on the predicted label count, encouraging the model to maximize rewards with fewer prediction words.
@@
 From $P_{1}(\cdot)$ to $P_{3}(\cdot)$, the penalty on prediction lengths becomes looser.
-In Section 5.3, we conduct experiments to evaluate the effects of different penalty strategies on emotion recognition performance and their efficacy in mitigating redundant predictions.
+In Section 5.3, we conduct experiments to evaluate the effects of different penalty strategies on recognition performance and their efficacy in mitigating redundant predictions.
 4 Experimental Database and Setup
 4.1 Corpus Description
-AffectGPT-R1 contains two training phases: cold-start training and reinforcement learning.
-In the first phase, we employ a large-scale dataset to equip the model with emotion understanding and format alignment capabilities.
-For this purpose, we select MER-Caption+ [lian2025affectgpt], which contains 31K automatically annotated emotion descriptions.
-In the second phase, we need a fine-grained dataset to better address OV-MER.
-Thus, we choose MER2025-OV [lian2025mer], which includes 1,200 samples with human-annotated open-vocabulary emotion labels.
-For performance comparison, we evaluate the model on OV-MERD+ [lian2025affectgpt] and MER-UniBench [lian2025affectgpt], assessing its capabilities in open-vocabulary and generalized emotion understanding.
-Appendix D provides additional details about MER-UniBench.
-Since 200 samples overlap between MER2025-OV and OV-MERD+, we exclude these samples from MER2025-OV to ensure a fair comparison, resulting in 1,000 samples used in reinforcement learning.
+AffectGPT-R1 consists of two training phases: cold-start training and reinforcement learning.
+In the first phase, we utilize a large-scale dataset to endow the model with emotion understanding and format alignment capabilities.
+For this purpose, we select MER-Caption+ (Lian et al., 2025a), which contains 31K samples annotated with rich emotion descriptions and open-vocabulary labels.
+In the second phase, we require a high-quality dataset specifically designed for OV-MER.
+Accordingly, we choose MER2025-OV (Lian et al., 2025b), which features human-annotated, high-quality open-vocabulary labels.
+For evaluation, we assess the model’s performance on OV-MERD+ (Lian et al., 2025a) and MER-UniBench (Lian et al., 2025a), measuring its capabilities in open-vocabulary and generalized emotion understanding.
+We conduct a rigorous overlap check to ensure there are no overlapping samples between the training and testing data.
+More details are provided in Appendix G.
 Table 1 presents statistics for these datasets.
@@
 In this table, “Sen.”, “Bas.”, “OV”, and “Des.” are abbreviations for sentiment labels, basic emotions, open-vocabulary emotions, and descriptive emotions, respectively.
+We rigorously verify that there is no overlap between the training and testing data.
+More details can be found in Appendix G.
 Dataset #Samples Task
 Sen. Bas. OV Des.
-Cold-start training
-MER-Caption+ [lian2025affectgpt] 31,327 $\times$ $\times$ $\surd$ $\surd$
-Reinforcement learning
-MER2025-OV [lian2025mer] 1,000 $\times$ $\times$ $\surd$ $\times$
+Training Phase 1: Cold-start training
+MER-Caption+ (Lian et al., 2025a) 31,327 $\times$ $\times$ $\surd$ $\surd$
+Training Phase 2: Reinforcement learning
+MER2025-OV (Lian et al., 2025b) 1,000 $\times$ $\times$ $\surd$ $\times$
 Testing
-OV-MERD+ [lian2025affectgpt] 532 $\times$ $\times$ $\surd$ $\times$
-MER-UniBench [lian2025affectgpt] 12,799 $\surd$ $\surd$ $\surd$ $\times$
+OV-MERD+ (Lian et al., 2025a) 532 $\times$ $\times$ $\surd$ $\times$
+MER-UniBench (Lian et al., 2025a) 12,799 $\surd$ $\surd$ $\surd$ $\times$
 4.2 Implementation Details
@@
 Due to memory constraints, we use a batch size of 3 for cold-start training and a batch size of 1 for reinforcement learning.
-For the model architecture, we adopt the default hyperparameters of AffectGPT [lian2025affectgpt].
+For the model architecture, we adopt the default hyperparameters of AffectGPT (Lian et al., 2025a).
 This setup ensures a fair comparison between AffectGPT and AffectGPT-R1, better highlighting the effectiveness of our proposed framework.
 5 Results and Discussion
-In this section, we first analyze the effectiveness of different reward functions, the impact of the output mode, and the effect of varying length penalties.
-Next, we compare the performance of AffectGPT-R1 with other baselines on MER-UniBench, demonstrating the effectiveness of our method for generalized emotion understanding.
-Finally, we conduct a series of ablation studies to investigate the influence of the choice of $G$, as well as the amount of data used in cold-start training and reinforcement learning.
 5.1 Reward Selection
@@
 Meanwhile, the format reward and alignment reward show limited impact, because these aspects are already well-learned during the cold-start phase.
-The weak performance of the perception reward may be due to current MLLMs’ inability to accurately decode human preferences, which could introduce reward noise [lian2025emoprefer].
+The weak performance of the perception reward may be due to current MLLMs’ inability to accurately decode human preferences, which could introduce reward noise (Lian et al., 2026a).
 Interestingly, the accuracy reward and dual reward yield similar performance, suggesting that while EW-based metrics are crucial, the choice of optimizing on thinking $o^{t}$ or answer $o^{a}$ has minimal impact.
@@
 Reward Functions OV-MERD+
-accuracy format alignment dual preception
+accuracy format alignment dual perception
 Without reinforcement learning
@@
 5.2 Output Mode Analysis
-In single-reward setups (see Table 2), the most effective rewards are the accuracy reward and dual reward, both of which are directly optimized on EW-based metrics.
-This observation leads to a heuristic hypothesis: Can we modify the output mode to generate and optimize solely on $o^{a}$, thereby making the model more focused on EW-based metrics?
-Specifically, for cold-start training, we remove the thinking process and require the output to contain only $o^{a}$.
-For reinforcement learning, we optimize the accuracy reward on $o^{a}$ directly.
-We denote this modified model as AffectGPT-R1 (w/o think) and the original model in Figure 1 as AffectGPT-R1 (w/ think).
-Experimental results in Table 3 show that AffectGPT-R1 (w/o think) achieves better performance.
-This improvement can be attributed to two reasons.
-Firstly, our thinking process is derived from coarse-grained datasets, where dataset noise may introduce incorrect reasoning and degrade performance.
-Secondly, including additional thinking content diverts the model’s attention from addressing OV-MER, thereby increasing training difficulty.
+In single-reward setups (see Table 2), the most effective rewards are the accuracy reward and dual reward, both of which are directly optimized based on EW metrics.
+This observation motivates a heuristic hypothesis: Could we modify the output mode to generate and optimize solely on $o^{a}$, thereby encouraging the model to focus more on EW-based metrics?
+Specifically, during cold-start training, we eliminate the thinking process and require the output to contain only $o^{a}$.
+During reinforcement learning, we optimize the accuracy reward directly on $o^{a}$.
+We refer to the modified model as AffectGPT-R1 (w/o think), and the original model shown in Figure 1 as AffectGPT-R1 (w/ think).
+In Table 3, we observe that removing the thinking process yields improved performance.
+This improvement can be attributed to two factors.
+First, our thinking process is derived from coarse-grained datasets, where inherent noise may lead to flawed reasoning and consequently degrade performance.
+Second, incorporating additional thinking content distracts the model from solving OV-MER, thereby increasing training difficulty.
+Regardless of whether the thinking process proves useful, this work represents a pioneering work in applying reinforcement learning to OV-MER.
+Future work will involve designing high-quality cold-start data and conducting an in-depth analysis to determine which samples or emotional states require additional thinking processes.
 Table 3: Output mode impact.
@@
 5.3 Impact of Different Penalties
-During inference, the model tends to generate lengthy responses with synonyms.
-This behavior primarily stems from the suboptimal design of the reward function.
+During inference, the model tends to generate lengthy predictions containing redundant emotion words.
+This behavior stems from the suboptimal design of the reward function.
 To mitigate this issue, we introduce three types of penalties in Section 3.3.3, where the penalty strength increases from $P_{3}(\cdot)$ to $P_{1}(\cdot)$.
@@
 Table 4: Impact of different penalties.
-The penalty strength increases from $P_{3}(\cdot)$ to $P_{1}(\cdot)$.
-Here, “overlong” indicates the presence of overlong outputs, and “degradation” indicates the presence of reward degradation.
-We seek a length penalty that preserves accuracy, prevents overlong outputs, and avoids causing reward degradation.
-Therefore, $P_{2}(\cdot)$ and $P_{3}(\cdot)$ are better choices.
-Penalty Accuracy Overlong Degradation
+Our goal is to identify a penalty that maintains model performance, prevents overlong outputs, and avoids introducing reward degradation.
+Based on these criteria, $P_{2}(\cdot)$ and $P_{3}(\cdot)$ are more suitable choices.
+Penalty Performance Overlong Degradation
 $P_{1}(\cdot)$ 67.44
@@
 None 68.39
-5.4 Main Results on MER-UniBench
-The above sections primarily focus on open-vocabulary emotions.
-In this section, we broaden the evaluation scope to MER-UniBench [lian2025affectgpt], a unified benchmark encompassing three typical tasks: sentiment analysis, basic emotion recognition, and fine-grained emotion detection.
-Experimental results are presented in Table 5.
-For fairness, all comparisons are conducted under the same input modality.
-These results demonstrate that AffectGPT-R1 achieves notable performance improvements in basic and fine-grained emotion recognition but exhibits slight degradation in sentiment analysis.
-This discrepancy arises because fine-grained and basic MER primarily focus on discrete emotion words (e.g., happy), whereas sentiment analysis centers on emotion polarity (e.g., positive and negative).
-Consequently, fine-grained and basic MER are more closely related to each other.
-Based on the average results across three tasks, AffectGPT-R1 achieves state-of-the-art performance on MER-UniBench, demonstrating the advantages of reinforcement learning for multimodal emotion understanding.
 Table 5: Main results on MER-UniBench.
@@
 Input Modality: Audio, Text
-OneLLM [han2024onellm] 64.01 54.09 63.39 61.98 25.52 17.21 28.32 33.44 22.25 41.14
-SECap [xu2024secap] 55.76 54.18 59.51 57.41 40.95 52.46 25.56 36.92 36.97 46.64
-PandaGPT [su2023pandagpt] 66.06 61.33 62.93 58.88 33.57 39.04 31.91 36.55 31.33 46.84
-Qwen-Audio [chu2023qwen] 70.09 46.90 70.73 65.26 41.85 31.61 49.09 35.47 32.36 49.26
-SALMONN [tang2023salmonn] 81.00 67.03 68.69 65.93 55.53 45.38 45.62 46.84 45.00 57.89
-AffectGPT [lian2025affectgpt] 83.46 80.74 82.99 83.75 72.94 73.41 56.63 55.68 59.98 72.18
-AffectGPT-R1 80.13 80.01 84.49 82.31 81.69 93.49 63.74 63.85 65.49 77.24
+OneLLM (Han et al., 2024) 64.01 54.09 63.39 61.98 25.52 17.21 28.32 33.44 22.25 41.14
+SECap (Xu et al., 2024) 55.76 54.18 59.51 57.41 40.95 52.46 25.56 36.92 36.97 46.64
+PandaGPT (Su et al., 2023) 66.06 61.33 62.93 58.88 33.57 39.04 31.91 36.55 31.33 46.84
+Qwen-Audio (Chu et al., 2023) 70.09 46.90 70.73 65.26 41.85 31.61 49.09 35.47 32.36 49.26
+SALMONN (Tang et al., 2023) 81.00 67.03 68.69 65.93 55.53 45.38 45.62 46.84 45.00 57.89
+AffectGPT (Lian et al., 2025a) 83.46 80.74 82.99 83.75 72.94 73.41 56.63 55.68 59.98 72.18
+AffectGPT-R1 (Ours) 80.13 80.01 84.49 82.31 81.69 93.49 63.74 63.85 65.49 77.24
 Input Modality: Video, Text
-Otter [li2025otter] 52.89 50.44 57.56 53.12 16.41 14.65 22.57 29.08 16.63 34.82
-Video-LLaVA [lin2024video] 56.37 61.64 53.28 57.45 36.93 30.25 30.73 38.95 34.00 44.40
-PandaGPT [su2023pandagpt] 58.50 64.25 62.07 65.25 39.13 47.16 38.33 47.21 35.07 50.77
-Video-ChatGPT [maaz2024video] 54.42 63.12 64.82 65.80 44.86 46.80 37.33 56.83 39.80 52.64
-VideoChat2 [li2024mvbench] 66.84 54.32 69.49 70.66 33.67 54.50 36.64 48.70 39.21 52.67
-LLaMA-VID [li2024llama] 61.78 63.89 69.35 67.48 50.72 57.60 42.75 46.02 45.01 56.07
-VideoChat [li2023videochat] 65.13 63.61 69.52 72.14 48.73 57.30 41.11 48.38 44.52 56.71
-Chat-UniVi [jin2024chat] 54.53 63.18 68.15 66.36 57.62 65.67 45.61 52.37 48.00 57.94
-mPLUG-Owl [ye2023mplug] 72.40 72.91 72.13 75.00 56.86 59.89 49.11 55.54 48.18 62.45
-AffectGPT [lian2025affectgpt] 82.39 81.57 87.20 86.29 74.58 75.29 57.63 62.19 61.65 74.31
-AffectGPT-R1 78.78 79.07 85.91 85.85 77.72 85.29 61.09 67.42 62.42 75.95
+Otter (Li et al., 2025a) 52.89 50.44 57.56 53.12 16.41 14.65 22.57 29.08 16.63 34.82
+Video-LLaVA (Lin et al., 2024) 56.37 61.64 53.28 57.45 36.93 30.25 30.73 38.95 34.00 44.40
+PandaGPT (Su et al., 2023) 58.50 64.25 62.07 65.25 39.13 47.16 38.33 47.21 35.07 50.77
+Video-ChatGPT (Maaz et al., 2024) 54.42 63.12 64.82 65.80 44.86 46.80 37.33 56.83 39.80 52.64
+VideoChat2 (Li et al., 2024a) 66.84 54.32 69.49 70.66 33.67 54.50 36.64 48.70 39.21 52.67
+LLaMA-VID (Li et al., 2024b) 61.78 63.89 69.35 67.48 50.72 57.60 42.75 46.02 45.01 56.07
+VideoChat (Li et al., 2025b) 65.13 63.61 69.52 72.14 48.73 57.30 41.11 48.38 44.52 56.71
+Chat-UniVi (Jin et al., 2024) 54.53 63.18 68.15 66.36 57.62 65.67 45.61 52.37 48.00 57.94
+mPLUG-Owl (Ye et al., 2023) 72.40 72.91 72.13 75.00 56.86 59.89 49.11 55.54 48.18 62.45
+AffectGPT (Lian et al., 2025a) 82.39 81.57 87.20 86.29 74.58 75.29 57.63 62.19 61.65 74.31
+AffectGPT-R1 (Ours) 78.78 79.07 85.91 85.85 77.72 85.29 61.09 67.42 62.42 75.95
 Input Modality: Audio, Video, Text
-PandaGPT [su2023pandagpt] 61.92 67.61 68.38 67.23 40.21 51.89 37.88 44.04 37.12 52.92
-R1-Omni [zhao2025r1] 58.02 56.48 71.82 68.58 64.17 67.43 43.20 51.58 55.24 59.61
-Emotion-LLaMA [cheng2024emotion] 66.13 67.66 78.32 77.23 59.38 73.62 46.76 55.47 52.97 64.17
-AffectGPT [lian2025affectgpt] 81.30 80.90 88.49 86.18 78.54 78.80 55.65 60.54 62.52 74.77
-AffectGPT-R1 79.65 80.18 87.26 85.75 84.51 93.13 66.71 74.26 68.39 79.98
-5.5 Ablation Study
-Figure 2: Left: Impact of $G$.
-“–” denotes the baseline model without reinforcement learning.
-Mid: Impact of data amount in cold-start training.
-This figure presents the performance of models trained with varying amounts of cold-start data and analyzes their influence on subsequent reinforcement learning.
-Right: Impact of data amount in reinforcement learning.
+PandaGPT (Su et al., 2023) 61.92 67.61 68.38 67.23 40.21 51.89 37.88 44.04 37.12 52.92
+R1-Omni (Zhao et al., 2025) 58.02 56.48 71.82 68.58 64.17 67.43 43.20 51.58 55.24 59.61
+Emotion-LLaMA (Cheng et al., 2024) 66.13 67.66 78.32 77.23 59.38 73.62 46.76 55.47 52.97 64.17
+AffectGPT (Lian et al., 2025a) 81.30 80.90 88.49 86.18 78.54 78.80 55.65 60.54 62.52 74.77
+AffectGPT-R1 (Ours) 79.39 79.24 88.25 84.97 84.32 93.75 63.12 74.26 68.05 79.48
+Figure 2: Left: Impact of data amount in cold-start training.
+This figure shows the performance of models trained with varying amounts of cold-start data and analyzes their influence on subsequent reinforcement learning.
+Mid: Impact of data amount in reinforcement learning.
 “–” denotes the baseline model without reinforcement learning.
 This figure shows that using a larger amount of data in reinforcement learning leads to better performance.
-Impact of $G$.
-GRPO requires sampling $G$ responses for group normalization (see Section 3.3.1).
-Figure 2 (left) reveals the impact of $G$.
-Due to GPU memory constraints, we limit our comparison to values of $G$ ranging from 2 to 16 and do not evaluate larger values.
-Experimental results demonstrate that, regardless of the chosen $G$, the model with reinforcement learning outperforms the one without reinforcement learning, demonstrating the effectiveness of the R1-style strategy in OV-MER.
-However, the selection of $G$ also matters.
-A small $G$ (e.g., $G=2$) fails to fully leverage the advantages of GRPO, primarily because insufficient sampling limits the likelihood of obtaining optimal results.
-Conversely, a larger $G$ (e.g., $G\geq 4$) ensures the effectiveness of GRPO, but further increasing $G$ (e.g., $G=16$) yields no significant improvement.
-Given that a larger $G$ also imposes higher GPU memory requirements, we adopt $G=4$ as our default setting.
+Right: Impact of $G$.
+“–” denotes the baseline model without reinforcement learning.
+5.4 Main Results on MER-UniBench
+Global-level Analysis.
+The preceding sections focus on open-vocabulary emotion recognition.
+In this section, we extend our evaluation to MER-UniBench (Lian et al., 2025a), a comprehensive benchmark that encompasses three tasks: sentiment analysis, basic emotion recognition, and fine-grained emotion detection.
+Appendix G provides additional details about this benchmark.
+Experimental results are summarized in Table 5.
+To ensure a fair comparison, all methods are evaluated using the same input modality.
+Based on the average performance across three tasks, AffectGPT-R1 achieves state-of-the-art results on MER-UniBench, demonstrating the effectiveness of reinforcement learning in enhancing emotion understanding.
+Task-level Analysis.
+Through further analysis of each task, AffectGPT-R1 demonstrates significant performance improvements in both basic and fine-grained emotion recognition, but shows a slight decline in sentiment analysis.
+This discrepancy stems from the fact that this work takes OV-MER as its primary focus and specifically tailors reward functions for this task.
+OV-MER aims to recognize discrete emotion words (e.g., happy and surprise), which closely align with the objectives of both basic and fine-grained emotion recognition.
+However, it differs somewhat from sentiment analysis, which targets emotional polarity (e.g., positive and negative).
+This distinction results in the reinforcement learning-optimized model, AffectGPT-R1, performing slightly worse on sentiment analysis compared to the non-optimized AffectGPT.
+Future work involves incorporating sentiment-related rewards into policy updates.
+5.5 Ablation Study
 Data Amount in Cold-start Training.
 AffectGPT-R1 consists of two stages: cold-start training (denoted as $s_{1}$) and reinforcement learning (denoted as $s_{2}$).
-In this section, we investigate the impact of the amount of cold-start data on $s_{1}$ and $s_{2}$.
-Experimental results in Figure 2 (mid) demonstrate that increasing the amount of cold-start data improves performance in $s_{1}$.
-Meanwhile, we observe that the performance after reinforcement learning is closely tied to the performance of the base model from $s_{1}$, and more cold-start data also enhances $s_{2}$ performance.
-The main reason lies in that reinforcement learning predominantly serves to reinforce correct answers rather than infuse novel knowledge [liu2025understanding].
-Therefore, a high-quality base model is essential for maximizing the effectiveness of reinforcement learning.
+In this section, we examine the impact of the amount of cold-start data on both $s_{1}$ and $s_{2}$.
+Experimental results shown in Figure 2 (left) indicate that increasing the amount of cold-start data improves performance in $s_{1}$.
+Furthermore, the performance after reinforcement learning strongly correlates with the performance of the base model from $s_{1}$, and additional cold-start data also benefits $s_{2}$.
+This is primarily because reinforcement learning mainly reinforces correct responses rather than introducing new knowledge (Liu et al., 2025).
+Therefore, a high-quality base model is crucial for maximizing the effectiveness of reinforcement learning.
 Data Amount in Reinforcement Learning.
-This section investigates the influence of data quantity on reinforcement training.
-Specifically, we start from the same model obtained through cold-start training (i.e., the model trained on 100% data in Figure 2 (mid)) and evaluate its performance with different amounts of reinforcement learning data.
-Experimental results in Figure 2 (right) demonstrate that even using just 10 samples for reinforcement learning leads to performance improvements, and further increasing the data amount yields more substantial gains.
+This section reveals the impact of data quantity on reinforcement learning.
+Specifically, starting from the same model obtained via cold-start training (i.e., the model trained on 100% data, as shown in the left part of Figure 2), we evaluate its performance using varying amounts of reinforcement learning data.
+Experimental results are shown in Figure 2 (mid).
+We observe that even using just 10 samples for reinforcement learning leads to performance improvements, confirming the effectiveness of reinforcement learning in addressing OV-MER.
+Moreover, increasing the amount of training data leads to greater performance gains.
 These phenomenon indicates that AffectGPT-R1 has not yet reached its upper-bound performance.
 In the future, we plan to collect more data to further improve the performance of AffectGPT-R1.
+Impact of $G$.
+GRPO requires sampling $G$ outputs for group normalization.
+Figure 2 (right) reveals the impact of $G$.
+Due to GPU memory limitations, we restrict our comparisons to values of $G$ between 2 and 16 and do not evaluate larger values.
+Experimental results show that, regardless of the specific value of $G$, the model trained with reinforcement learning consistently outperforms the one without it, confirming the effectiveness of our approach.
+Meanwhile, the selection of $G$ also matters.
+A small $G$ does not fully exploit the benefits of GRPO, mainly because limited sampling reduces the chances of identifying optimal outcomes.
+In contrast, a larger $G$ (e.g., $G\geq 4$) ensures the effectiveness of GRPO, but further increasing $G$ yields no significant improvement.
+Since a larger $G$ also increases computational cost, we set $G=4$ as the default configuration.
 6 Conclusion
-This paper proposes AffectGPT-R1, the first work applying reinforcement learning to OV-MER.
+This paper proposes AffectGPT-R1, a pioneering work that applies reinforcement learning to OV-MER.
 We introduce various rewards and incorporate length penalties to mitigate reward hacking.
 Through extensive experiments, we reveal the impact of different rewards and identify the optimal reward design.
-AffectGPT-R1 demonstrates strong effectiveness in OV-MER, and this framework also achieves state-of-the-art performance on MER-UniBench, a benchmark for generalized emotion understanding.
-Furthermore, we examine the influence of various modules and hyperparameters, providing valuable insights for future research.
-By proving the potential of reinforcement learning as a promising approach for emotion understanding, our work will advance the development of this field toward a reinforcement learning era.
-[…]
-Supplementary Material
-A EW-based Metric
-In this section, we present a detailed calculation formula for EW-based metrics.
+AffectGPT-R1 shows strong effectiveness in OV-MER and also achieves state-of-the-art performance on MER-UniBench, a benchmark designed for generalized emotion understanding.
+Furthermore, we investigate the effects of various modules and hyperparameters, offering valuable insights for future research.
+By demonstrating the potential of reinforcement learning as a powerful paradigm for emotion understanding, our work will advance the development of this field toward a reinforcement learning era.
+Impact Statement
+Social Impact.
+Emotion plays an important role in mental health, human-computer interaction, and embodied AI. This paper aims to advance the development of MER techniques, thereby promoting their applications.
+Ethics Statement.
+We do not collect or annotate new data, but use existing datasets to verify the effectiveness of reinforcement learning in OV-MER.
+The use of all datasets has been authorized by their owners under corresponding usage requirements.
+Therefore, no ethical issues are raised.
+Appendix A Reproducibility Statement
+In this paper, we have made every effort to ensure the reproducibility of our work.
+In the supplementary material, we provide our source code and command lines for cold-start training, reinforcement learning, and inference.
+Additionally, we include a detailed README file.
+To comply with ICML’s anonymity requirements, we have removed all external links and owner names.
+After acceptance, we will release the full code on GitHub to facilitate further research.
+Appendix B Novelty Statement
+In this paper, we present pioneering work that reveals the potential of reinforcement learning in OV-MER.
+There are currently two main types of emotion recognition tasks: discriminative emotions and generative emotions.
+The former primarily involves tasks with a constrained label space, such as basic or dimensional emotions (Wang et al., 2023; Sun et al., 2024; Fang et al., 2025).
+In contrast, the latter imposes no restrictions on the label space, giving rise to the emerging area of open-vocabulary emotions (Lian et al., 2025c).
+Recent research has shown a growing interest in shifting from discriminative emotions to generative emotions, driven by the need for more nuanced emotional representations (Lian et al., 2025a; Cheng et al., 2024; Zhang et al., 2025a).
+However, effectively addressing generative emotions remains an open challenge.
+In this work, we focus on open-vocabulary emotions, a specific subtask within generative emotions, and propose leveraging reinforcement learning as a solution approach.
+During our investigation, several practical issues emerged, including the design of reward functions and strategies to mitigate reward hacking.
+By tackling these challenges, our experimental results demonstrate the effectiveness of reinforcement learning in handling generative emotions.
+Therefore, this paper should not be viewed merely as an application of reinforcement learning, but rather as a significant milestone in advancing generative emotions research.
+This work will inspire more researchers to consider reinforcement learning as a foundational tool for addressing generative emotions.
+Meanwhile, we would like to emphasize that this work differs from R1-Omni (Zhao et al., 2025), another RL-based framework in affective computing.
+Specifically, R1-Omni focuses on discriminative emotions, whereas AffectGPT-R1 shifts the task to generative emotions.
+Due to this task difference, we adopt distinct reward functions and face unique challenges, such as reward hacking, that are not present in R1-Omni.
+Meanwhile, it is important to note that R1-Omni is only available as a preprint on arXiv.
+Following the ICML policy, we are not required to compare with it.
+Nevertheless, to thoroughly validate the effectiveness of AffectGPT-R1, we include a comparison with R1-Omni, and the experimental results demonstrate our model’s advantages in open-vocabulary emotion detection and generalized emotion understanding.
+Appendix C Emotion Wheel (EW)
+The emotion wheel is a structured visual tool designed to categorize and identify human emotions.
+It organizes emotions into core categories (the inner sections of the wheel) and their nuanced variations (the outer sections), providing intuitive insights into emotional structure.
+This paper uses five emotion wheels to compute EW-based metrics, aligning with prior work (Lian et al., 2025a) to ensure fair comparisons.
+Figure 3 illustrates these emotion wheels.
+(a) W1
+(b) W2
+(c) W3
+(d) W4
+(e) W5
+Figure 3: Emotion wheels. This paper uses five emotion wheels to calculate EW-based metrics, ensuring a fair comparison with previous work (Lian et al., 2025a, c).
+Appendix D EW-based Metric
+In this section, we present a detailed calculation formula for EW-based metrics (Lian et al., 2025c).
 The computation method primarily consists of two components: eliminating the influence of synonyms and defining set-level metrics.
-A.1 Handling Synonyms
+D.1 Handling Synonyms
 To mitigate the impact of synonyms, we employ a three-level hierarchical grouping strategy:
-(L1) We map different forms of words to their base form. For example, happier and happiness are mapped to happy. This function is denoted as $F_{l_{1}}(\cdot)$.
-(L2) We map synonyms to a unified word. For example, happy and joyful are mapped to happy. This mapping function is denoted as $F_{l_{2}}(\cdot)$.
-(L3) Emotion wheel provides natural grouping information [plutchik1980general].
-Since there is no consensus on the emotion wheel, we adopt $K=5$ emotion wheels, consistent with prior work [lian2025ov].
-Appendix B visualizes all emotion wheels.
+(L1) We map different forms of words to their base form.
+For example, happier and happiness are mapped to happy.
+To achieve this, for a base word $w$, we first use GPT to generate various forms of $w$, with the output denoted as $l_{1}(w)$.
+Since different results may be produced across multiple calls, we repeat this process ten times and merge all the generated words, resulting in $\hat{l}_{1}(w)$.
+To map these different word forms back to their base form, we simply map $\hat{l}_{1}(w)$ to $w$, and we refer to this mapping function as $F_{l_{1}}(\cdot)$.
+(L2) We map synonyms to a unified word.
+For example, happy and joyful are mapped to happy.
+To achieve this, similar to L1, we use GPT to generate synonyms for a given word $w$, repeat the process multiple times, and then merge all the generated words, obtaining $\hat{l}_{2}(w)$.
+To map the synonyms to their unified form, we map each element in $\hat{l}_{2}(w)$ to $w$, and we refer to this mapping function as $F_{l_{2}}(\cdot)$.
+(L3) The emotion wheel offers natural grouping information for emotions (Plutchik, 1980).
+Since there is no consensus on the emotion wheel, we adopt $K=5$ distinct emotion wheels, consistent with prior work (Lian et al., 2025a, c) to ensure fair comparisons.
+Figure 3 visualizes all five emotion wheels.
 For each wheel $w_{k}$, we map all outer labels to their corresponding inner labels.
 This function is denoted as $F_{l_{3}}^{w_{k}}(\cdot)$.
-The clustering functions can be summarized as:
+Finally, the clustering functions can be summarized as:
 $G_{w_{k}}(\cdot)=F_{l_{3}}^{w_{k}}{\left(F_{l_{2}}\left(F_{l_{1}}\left(\cdot\right)\right)\right)},k\in[1,K].$ (16)
-A.2 Set-level Metric
+D.2 Set-level Metric
 Since we do not restrict the number of predicted and annotated emotions, we employ set-level evaluation metrics.
@@
 For a sample $x_{i}$, the true labels are $\mathbf{Y}_{i}=\{y_{i}^{j}\}_{j=1}^{n_{i}}$ and the predicted labels are $\mathbf{\hat{Y}}_{i}=\{\hat{y}_{i}^{j}\}_{j=1}^{\hat{n}_{i}}$.
+Here, $n_{i}$ and $\hat{n}_{i}$ are the number of predicted and annotated emotions, respectively.
 The evaluation metrics are defined as follows:
@@
 $\text{EW}\left(\mathbf{Y}_{i},\mathbf{\hat{Y}}_{i}\right)=\frac{1}{K}\sum_{k=1}^{K}\mbox{F}_{\mbox{s}}^{k}.$ (20)
-B Emotion Wheel
-The emotion wheel is a structured visual tool designed to categorize and identify human emotions.
-It organizes emotions into core categories (the inner sections of the wheel) and their nuanced variations (the outer sections), providing intuitive insights into emotional structure.
-This paper uses five emotion wheels to compute metrics for OV-MER, aligning with prior work [lian2025affectgpt] to ensure fair comparisons.
-(a) W1
-(b) W2
-(c) W3
-(d) W4
-(e) W5
-Figure 3: Emotion wheels. This paper uses five emotion wheels to calculate EW-based metrics.
-C Baseline Details
-SECap [xu2024secap].
+Appendix E Details of Alignment Reward
+E.1 Emotion Word Extraction
+To compute the alignment reward, we need to extract emotion words $e^{t}$ from $o^{t}$.
+To do so, we input $o^{t}$ into Qwen2.5-7B using the following prompt: Please assume the role of an expert in the field of emotions.
+We provide clues related to the emotional state of a character.
+Based on the provided clues, please identify the character’s emotional states.
+Separate different emotional categories with commas, and output only the clearly identifiable emotional categories in a list format.
+If no emotions can be identified, please return an empty list.
+E.2 Similarity Calculation
+Alignment reward involves computing the similarity between two emotion sets.
+In practice, we use the Jaccard Similarity Coefficient, which measures the similarity between two sets by comparing the size of their intersection to the size of their union.
+For clarity, we adopt the notations defined in Appendix D, and the method for calculating the Jaccard Similarity Coefficient is provided below:
+$\mbox{Similarity}_{\mbox{s}}^{k}=\frac{1}{N}\sum_{i=1}^{N}\frac{\left|G_{w_{k}}(\mathbf{Y}_{i})\cap G_{w_{k}}(\mathbf{\hat{Y}}_{i})\right|}{\left|G_{w_{k}}(\mathbf{Y}_{i})\cup G_{w_{k}}(\mathbf{\hat{Y}}_{i})\right|},$ (21)
+$\mathrm{is\_similar}\left(\mathbf{Y}_{i},\mathbf{\hat{Y}}_{i}\right)=\frac{1}{K}\sum_{k=1}^{K}\mbox{Similarity}_{\mbox{s}}^{k}.$ (22)
+Appendix F Baseline Details
+SECap (Xu et al., 2024).
 This is a speech emotion captioning framework designed to describe speech emotions using natural language.
@@
 Finally, these features are fed into the LLM, which generates natural language captions describing speech emotions.
-SALMONN [tang2023salmonn].
+SALMONN (Tang et al., 2023).
 This model is capable of processing both speech and audio.
-Specifically, it employs Whisper [radford2023robust] to encode speech and BEATs [chen2023beats] to encode audio.
+Specifically, it employs Whisper (Radford et al., 2023) to encode speech and BEATs (Chen et al., 2023b) to encode audio.
 The resulting features are then integrated into the LLM via a window-level Q-Former.
 Additionally, a LoRA adapter is applied to enhance the LLM’s instruction-following and question-answering capabilities.
-Qwen-Audio [chu2023qwen].
+Qwen-Audio (Chu et al., 2023).
 This model is capable of handling 30 different tasks and a variety of audio types, demonstrating universal audio understanding capabilities.
@@
 The output features are then passed to LLMs for answer generation.
-Otter [li2025otter].
-This model is built upon the OpenFlamingo framework [awadalla2023openflamingo] and trained on the MIMIC-IT dataset [li2025otter], demonstrating strong capabilities in multimodal perception, reasoning, and in-context learning.
-Chat-UniVi [jin2024chat].
+Otter (Li et al., 2025a).
+This model is built upon the OpenFlamingo framework (Awadalla et al., 2023) and trained on the MIMIC-IT dataset (Li et al., 2025a), demonstrating strong capabilities in multimodal perception, reasoning, and in-context learning.
+Chat-UniVi (Jin et al., 2024).
 This is a unified vision-language model capable of processing both images and videos. It employs dynamic visual tokens that capture spatial details in images and temporal relationships in videos.
-Video-LLaVA [lin2024video].
+Video-LLaVA (Lin et al., 2024).
 It aligns images and videos before projection, enabling the LLM to learn from a unified visual representation and equipping it with the ability to comprehend both images and videos simultaneously.
-Video-ChatGPT [maaz2024video].
+Video-ChatGPT (Maaz et al., 2024).
 This model averages frame-level features across temporal and spatial dimensions, extracting both spatial and temporal video features.
 These features are then projected into the text token space, enabling the LLM to generate responses.
-VideoChat [li2023videochat].
+VideoChat (Li et al., 2025b).
 It uses video foundation models to encode videos as embeddings, then employs LLMs for response generation.
 After instruction fine-tuning, the model demonstrates promising capabilities across various video applications.
-VideoChat2 [li2024mvbench].
+VideoChat2 (Li et al., 2024a).
 This model leverages diverse instruction-tuning data and employs a three-stage progressive multimodal training approach, comprising vision-language alignment, vision-language connection, and instruction tuning.
 As a result, it outperforms leading models across multiple tasks.
-LLaMA-VID [li2024llama].
+LLaMA-VID (Li et al., 2024b).
 This model represents each image with two tokens: a context token and a content token.
@@
 By condensing each image into just a few tokens, the model can efficiently understand and process long videos.
-mPLUG-Owl [ye2023mplug].
+mPLUG-Owl (Ye et al., 2023).
 This approach equips LLMs with vision-language capabilities through a two-stage training paradigm, enabling strong performance on instruction-following, visual understanding, and reasoning tasks.
-PandaGPT [su2023pandagpt].
-It uses ImageBind [girdhar2023imagebind] to map multiple modalities into a shared embedding space, then employs a projector to transform these features into the LLM’s input space, and finally leverages the LLM to generate responses.
-OneLLM [han2024onellm].
+PandaGPT (Su et al., 2023).
+It uses ImageBind (Girdhar et al., 2023) to map multiple modalities into a shared embedding space, then employs a projector to transform these features into the LLM’s input space, and finally leverages the LLM to generate responses.
+OneLLM (Han et al., 2024).
 This model aligns eight modalities with language through a unified framework.
@@
 Additionally, it employs a progressive multimodal alignment pipeline, enabling it to comprehend eight input types and generate responses based on user queries.
-Emotion-LLaMA [cheng2024emotion].
+R1-Omni (Zhao et al., 2025).
+This framework also leverages reinforcement learning for emotion recognition.
+Unlike AffectGPT-R1, which focuses on open-vocabulary emotions, this work targets basic emotions.
+Due to the different objectives, the reward functions used in R1-Omni differ from those employed in AffectGPT-R1.
+A more detailed comparison between R1-Omni and AffectGPT-R1 is provided in Appendix B.
+Emotion-LLaMA (Cheng et al., 2024).
 This framework follows the architecture of traditional MLLMs and consists of three key components: encoders, projectors, and LLMs.
@@
 Experimental results on benchmark datasets demonstrate that this framework achieves improved performance over previous solutions.
-R1-Omni [zhao2025r1].
-It focuses on basic emotion recognition and uses reinforcement learning with verifiable rewards to address this task.
-Unlike R1-Omni, AffectGPT-R1 shifts the task to open-vocabulary emotion detection.
-Due to this task difference, we adopt distinct reward functions and face unique challenges, such as reward hacking, that are not present in R1-Omni.
-It is important to note that R1-Omni is only available as a preprint on arXiv.
-Following CVPR policy, we are not required to compare with it.
-Nevertheless, to thoroughly validate the effectiveness of AffectGPT-R1, we include a comparison with R1-Omni, and the experimental results demonstrate our model’s advantages in open-vocabulary emotion detection and generalized emotion understanding.
-Additionally, unlike R1-Omni, which only releases its inference code, we provide both training and inference code to enhance reproducibility.
-AffectGPT [lian2025affectgpt].
+AffectGPT (Lian et al., 2025a).
 This framework differs from previous MLLMs, which typically delegate the entire multimodal fusion process to the language model.
@@
 Experimental results demonstrate the effectiveness of AffectGPT in generalized emotion understanding, including sentiment analysis, basic emotion recognition, and fine-grained emotion detection.
-D MER-UniBench Details
-MER-UniBench is a unified benchmark that covers three typical tasks, including sentiment analysis, basic emotion recognition, and fine-grained emotion detection, with the aim of evaluating the generalized emotion understanding capabilities of different models.
-In Table 6, we summarize the data components for each task.
-Please refer to the original paper for more details [lian2025affectgpt].
-Table 6: MER-UniBench. This benchmark encompasses three tasks, and this table presents the data components for each task. In total, this benchmark includes 12,799 samples for testing.
-Raw Dataset Selected Subset # Samples
+Appendix G Dataset Details
+G.1 Training Datasets
+MER-Caption+ (Lian et al., 2025a).
+MER-Caption+ is a large-scale descriptive emotion dataset comprising approximately 31K samples and 2K emotion labels.
+To reduce annotation costs, it introduces a novel model-led human-assisted strategy that leverages human priors to guide both description generation and sample filtering.
+Specifically, in the description generation stage, the pre-trained models are carefully selected to maximize performance on fine-grained emotion detection.
+During sample filtering, a two-level filtering process is employed to remove samples with mismatched audio and video, abnormally long descriptions, and those that perform poorly on sentiment analysis and basic emotion recognition.
+MER2025-OV (Lian et al., 2025b).
+MER2025-OV is specifically designed for OV-MER.
+Unlike the automatically annotated MER-Caption+, MER2025-OV is a high-quality dataset created through a purely manual annotation process.
+It employs a multi-round annotation strategy.
+Initially, a pool of candidate emotions is provided, and four experts in affective computing are asked to select the correct labels and add any missing ones.
+All selected labels are combined to form the candidate set for the next round.
+In the second round, four annotators perform another round of labeling, and only labels selected by at least two annotators are retained.
+This two-stage approach ensures both completeness, by accepting labels chosen by any annotator in the first round, and accuracy, by filtering for consensus in the second.
+Through this rigorous process, MER2025-OV guarantees the high quality of its annotated emotion labels.
+G.2 Testing Datasets
+OV-MERD+ (Lian et al., 2025a).
+OV-MERD+ is a dataset specifically designed for OV-MER, which is an extended version of OV-MERD (Lian et al., 2025c).
+The original OV-MERD dataset consists of 332 samples, all of which are fully annotated by humans for open-vocabulary emotions.
+Following a similar annotation process, OV-MERD+ expands the dataset to a total of 532 samples.
+In this study, we employ OV-MERD+ as the test set to evaluate the performance of different models on open-vocabulary emotion detection.
+MER-UniBench (Lian et al., 2025a).
+MER-UniBench is a benchmark that encompasses three representative tasks: sentiment analysis, basic emotion recognition, and fine-grained emotion detection.
+Specifically, sentiment analysis focuses on binary classification into positive and negative sentiments.
+Basic emotion recognition restricts the label space to a predefined set of basic emotion categories.
+In contrast, fine-grained emotion detection removes these constraints, allowing predictions across any category.
+This benchmark is designed to assess generalized emotion understanding capabilities.
+Table 6 summarizes the data composition for each task.
+Table 6: MER-UniBench.
+This table presents the data components for three tasks: sentiment analysis, basic emotion recognition, and fine-grained emotion detection.
+In total, this benchmark includes 12,799 test samples.
+Raw Dataset Selected Subset # Samples Label Space
 Sentiment Analysis
-CMU-MOSI [zadeh2017tensor] Test 686
-CMU-MOSEI [zadeh2018multimodal] Test 4,659
-CH-SIMS [yu2020ch] Test 457
-CH-SIMS v2 [liu2022make] Test 1,034
+CMU-MOSI (Zadeh et al., 2017) Test 686 positive, negative
+CMU-MOSEI (Zadeh et al., 2018) Test 4,659 positive, negative
+CH-SIMS (Yu et al., 2020) Test 457 positive, negative
+CH-SIMS v2 (Liu et al., 2022) Test 1,034 positive, negative
 Basic Emotion Recognition
-MER2023 [lian2023mer] MER-MULTI 411
-MER2024 [lian2024mer] MER-SEMI 1,169
-IEMOCAP [busso2008iemocap] Session5 1,241
-MELD [poria2019meld] Test 2,610
+MER2023 (Lian et al., 2023a) MER-MULTI 411 worry, happy, neutral, angry, surprised, sad
+MER2024 (Lian et al., 2024a) MER-SEMI 1,169 worry, happy, neutral, angry, surprised, sad
+IEMOCAP (Busso et al., 2008) Session5 1,241 anger, happiness, sadness, neutral
+MELD (Poria et al., 2019a) Test 2,610 anger, joy, sadness, neutral, disgust, fear, surprise
 Fine-grained Emotion Detection
-OV-MERD+ [lian2025affectgpt] All 532
-E Impact Statement
-E.1 Social Impact
-Emotion plays an important role in mental health, human-computer interaction, and embodied AI.
-This paper aims to advance the development of MER techniques, thereby promoting their applications and improving our daily lives.
-E.2 Ethics Statement
-This paper does not collect or annotate new data but utilizes existing datasets to verify the effectiveness of reinforcement learning in OV-MER.
-The use of all datasets has been authorized by their owners under corresponding usage requirements.
-Therefore, no ethical issues are raised in this paper.
-F Reproducibility Statement
-In this paper, we have made every effort to ensure the reproducibility of our work.
-In the supplementary material, we provide our source code and command lines for cold-start training, reinforcement learning, and inference.
-Additionally, we include a detailed README file.
-To comply with CVPR’s anonymity requirements, we have removed all external links and owner names.
-After acceptance, we will release the full code on GitHub to facilitate further research.
+OV-MERD+ (Lian et al., 2025a) All 532 unrestricted label space
+G.3 Data Leakage Prevention and Overlap Removal
+To prevent data leakage, we perform a strict check to ensure that there is no overlap between the training and testing data.
+During this process, we identify 200 overlapping samples between MER2025-OV (part of the training set) and OV-MERD+ (part of the test set).
+Consequently, we remove these overlapping samples from MER2025-OV, resulting in 1,000 samples available for training (see Table 1).
+Apart from this case, we do not observe any other instances of overlap between the training and testing data.
+G.4 Task Clarification
+In this paper, we focus on utterance-level emotion recognition, as opposed to dialogue-level emotion recognition.
+The key difference between these tasks lies in whether conversational history is taken into account.
+For utterance-level emotion recognition, the input consists of a single video clip centered on one character (Tsai et al., 2019).
+In contrast, dialogue-level emotion recognition requires leveraging additional dialogue information to perform emotion inference (Poria et al., 2017).
+This distinction has led to the development of two types of benchmarks: the utterance-level benchmark (Lian et al., 2026b) and the dialogue-level benchmark (Poria et al., 2019b).
+In MER-UniBench, although certain datasets (e.g., MELD and IEMOCAP) contain dialogue information, others (e.g., MER2023 and CH-SIMS) do not.
+Furthermore, even when dialogue context is available in datasets like MELD and IEMOCAP, its inclusion is not strictly necessary and depends on the specific experimental setup (Tsai et al., 2019; Poria et al., 2017).
+Consequently, MER-UniBench is designed as an utterance-level benchmark, and this work adopts the same setting to ensure fair comparison with prior studies (Lian et al., 2025a).
+Dialogue-level emotion recognition is beyond the scope of this paper and will be explored in our future work.